### v1

# Implicit Actor Critic Coupling via a Supervised Learning Framework for RLVR

Abstract
Recent advances in Reinforcement Learning with Verifiable Rewards (RLVR) have empowered large language models (LLMs) to tackle challenging reasoning tasks such as mathematics and programming. RLVR leverages verifiable outcome rewards to guide policy optimization, enabling LLMs to progressively improve output quality in a grounded and reliable manner. Despite its promise, the RLVR paradigm poses significant challenges, as existing methods often suffer from sparse reward signals and unstable policy gradient updates, particularly in RL-based approaches. To address the challenges, we propose PACS, a novel RLVR framework that achieves imPlicit Actor Critic coupling via a Supervised learning framework. By treating the outcome reward as a predictable label, we reformulate the RLVR problem into a supervised learning task over a score function parameterized by the policy model and optimized using cross-entropy loss. A detailed gradient analysis shows that this supervised formulation inherently recovers the classical policy gradient update while implicitly coupling actor and critic roles, yielding more stable and efficient training. Benchmarking on challenging mathematical reasoning tasks, PACS outperforms strong RLVR baselines, such as PPO and GRPO, achieving superior reasoning performance. For instance, PACS achieves 59.78% at pass@256 on AIME 2025, representing improvements of 13.32 and 14.36 points over PPO and GRPO. This simple yet powerful framework offers a promising avenue for LLMs post-training with verifiable rewards. Our code and data are available as open source at https://github.com/ritzz-ai/PACS.
1 Introduction
Large Language Models (LLMs) have recently achieved remarkable success in complex reasoning tasks, driven by advancements, such as OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025). In particular, scaling test-time compute enables longer Chains of Thought (CoT), facilitating more sophisticated and structured reasoning. This progress has led to significant improvements in challenging domains, such as mathematics (Shao et al., 2024; Liu et al., 2025a; Zuo et al., 2025) and programming (Lambert et al., 2024; Cui et al., 2025; Zhao et al., 2025). A crucial underlying ingredient behind these achievements is the emergence of Reinforcement Learning with Verifiable Rewards (RLVR), which empowers LLMs to progressively enhance their capabilities by reinforcing their self-generated outputs with verifiable outcome rewards.
Existing RLVR methods fall into two primary categories: value-model-based approaches, such as PPO (Schulman et al., 2017) and VAPO (Yue et al., 2025), which employ large value models to estimate advantages for policy updates; and value-model-free approaches, such as GRPO (Shao et al., 2024; Guo et al., 2025) and DAPO (Yu et al., 2025), which bypass explicit value modeling by leveraging Monte Carlo estimation of group relative rewards. While both classes have demonstrated remarkable success, they face fundamental challenges stemming from the sparse nature of outcome rewards in RLVR settings, where only a single reward signal is provided after the entire response is generated. This sparsity severely hinders credit assignment and gradient propagation in Reinforcement Learning (RL)-based policy learning, which relies on attributing feedback to individual tokens or actions within the output. Value-model-based methods address this challenge by learning explicit value models, but at the cost of increased model complexity and computational overhead. In contrast, value-model-free methods avoid this complexity by forgoing explicit value prediction, but suffer from high variance in their Monte Carlo estimates, often leading to issues such as advantage collapse, training instability, and degraded performance.
Figure 1: Comparison between RLVR and the supervised learning reformulation, where the query and output are input, and the outcome reward is treated as a predictable label.
This challenge, rooted in the inherently sparse nature of RLVR feedback, highlights fundamental limitations of the existing RL-based paradigm. It motivates the development of alternative policy optimization strategies that can leverage direct supervision to enable more stable and efficient learning in RLVR settings. To this end, we propose a novel framework that recasts the RLVR problem as a supervised learning task (see Figure 1). Instead of relying on sparse outcome rewards to optimize the policy via RL, we directly interpret the outcome reward as a supervision signal and train a score function, parameterized by the policy itself, to predict this reward via a cross-entropy loss.
Through a detailed gradient analysis, we show that this supervised learning formulation not only recovers the standard policy gradient update but also implicitly couples the actor (policy improvement) and critic (reward estimation) components within a single update. Unlike conventional actor critic methods that maintain a separate value function estimator, our approach adopts a shared parameterization, allowing both roles to be updated jointly via a unified learning signal. This tight coupling eliminates the temporal mismatch between reward estimation and policy updates, thereby improving training efficiency. Moreover, in contrast to value-model-free methods, which rely solely on high-variance Monte Carlo estimates of relative advantages for policy updates, our method leverages prediction-error-based learning signals, providing more stable and reliable guidance for optimization. We refer to our method as PACS (imPlicit Actor Critic coupling via a Supervised learning framework). It offers a principled and efficient training paradigm that unifies policy learning and reward estimation within a coherent supervised learning framework.
To assess the effectiveness of our method, we conduct experiments on four challenging mathematical reasoning tasks and compare it against the state-of-the-art RLVR methods, including PPO and GRPO. The results demonstrate that PACS significantly outperforms baselines. For instance, PACS achieves a pass@256 rate of 59.78% on AIME 2025, representing improvements of 13.32 and 14.36 points over PPO and GRPO. Ablation studies confirm the critical importance of the weighting mechanism, while training dynamics analysis reveals that PACS achieves superior exploration-exploitation balance.
The main contributions are summarized as follows:
  • We propose PACS, a novel RLVR approach that reframes the problem as supervised learning: motivated by the sparsity of outcome rewards, we treat them as target labels and train the policy via cross-entropy loss.
  • We demonstrate that the proposed loss inherently captures policy gradient updates and implicitly unifies actor (policy improvement) and critic (reward estimation) through shared parameterization, enabling efficient and stable policy optimization.
  • Extensive experiments on challenging mathematical reasoning benchmarks show that PACS outperforms state-of-the-art RLVR methods, achieving superior reasoning performance, maintaining healthy policy entropy for exploration, and exhibiting improved training efficiency.
2 Related Work
Reasoning Models.
The emergence of reasoning models, referring to LLMs explicitly trained or optimized for multi-step chain-of-thought inference, has been catalyzed by the release of OpenAI-o1 (Jaech et al., 2024), which achieves strong performance on challenging reasoning tasks. DeepSeek-R1 (Guo et al., 2025) further highlights the power of verifiable reward signals by using GRPO (Shao et al., 2024) with rule-based outcome rewards to guide effective policy optimization. Building on this foundation, QWQ (Qwen-Team, 2025) introduces a two-stage RL framework: a “cold-start” phase to bootstrap reasoning ability, followed by a brief phase of general RL finetuning to expand the model’s problem-solving proficiency. Most recently, Gemini 2.5 Pro (Comanici et al., 2025) integrates human feedback, retrieval-augmented generation, and multimodal perception, enabling reasoning grounded in diverse information sources beyond text. Together, these advances illustrate the evolution of RL-enhanced reasoning models from single-domain systems to general-purpose, multimodal agents.
Reinforcement Learning with Verifiable Reward (RLVR).
RLVR has become a widely adopted strategy for enhancing the reasoning capabilities of LLMs in domains with clearly verifiable correctness, such as mathematics and programming  (Guo et al., 2025; Team et al., 2025; Ma et al., 2025; Chu et al., 2025). Recent work has explored a wide range of RLVR techniques, which can be broadly categorized as value-model-based or value-model-free. Value-model-based methods, such as PPO (Schulman et al., 2017), VinePPO (Kazemnejad et al., 2024), and VAPO (Yue et al., 2025), explicitly learn a value function to estimate the expected cumulative reward, providing stable training signals at the cost of additional computational overhead. On the other hand, value-model-free methods, such as GRPO (Shao et al., 2024), REINFORCE++ (Hu et al., 2025), and DAPO (Yu et al., 2025), avoid explicit value modeling by relying on Monte Carlo-based advantage estimation. While these approaches reduce modeling complexity, they often suffer from high gradient variance, which undermines training stability and overall performance.
3 Methodology
In this section, we first establish the preliminaries for RLVR, and then introduce PACS, a novel RLVR framework that achieves imPlicit Actor–Critic coupling through a Supervised learning framework. We present the overall framework, followed by a detailed gradient analysis that illustrates the effectiveness of our approach, and conclude with practical design considerations to instantiate the implementation.
3.1 Preliminaries
Let $\theta$ denote the parameters of an LLM, and let $P(Q)$ be a distribution over queries from which a query $q$ is sampled. The model generates an output $o$ according to the policy $\pi_{\theta}(\cdot|q)$. A verifiable outcome reward $R(q,o)\in\{0,1\}$ is provided by either a reward model or an external verifier which determines whether the model’s output $o$ to the query $q$ is correct ($1$) or incorrect (0). The objective of RLVR is to learn a policy $\pi_{\theta}$ that maximizes the expected reward: $\mathcal{L}_{\text{RLVR}}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(\cdot|q)}[R(q,o)]$. Various RLVR approaches (Schulman et al., 2017; Kool et al., 2019; Ahmadian et al., 2024; Shao et al., 2024; Yu et al., 2025; Yue et al., 2025) have been proposed to optimize this objective function, each seeking to effectively learn the policy $\pi_{\theta}$ under the supervision of outcome rewards.
3.2 Recasting RLVR within a Supervised Learning Framework
In the RLVR framework, LLMs generate responses to input queries, and verifiable outcome rewards that indicate the correctness of the final outcomes are provided at the end of responses. A key characteristic of this setting is that the reward is delivered only once, after the entire response is generated, reflecting the overall quality of the output. Inspired by this observation, we propose an alternative to conventional RL-based optimization: rather than policy learning through RL from sparse reward signals, we recast the problem as a supervised learning task in which the model is trained to directly predict the outcome reward.
Figure 2: An illustration of the PACS framework. The framework consists of three main components: (1) Reward Proxy Computation, which calculates a reward proxy $\hat{r}$ based on the log-probability ratio. (2) Group Computation, which computes RLOO-based advantage scores $\psi$ from the reward proxies. (3) Cross-Entropy Loss, which converts the RLVR problem into a supervised learning task, optimizing a scoring function parameterized by the policy with a cross-entropy loss.
Concretely, instead of using the reward $R(q,o)\in\{0,1\}$ associated with a query-output pair $(q,o)$ to guide policy optimization via RL, we consider $(q,o)$ as input data and treat $R(q,o)$ as the corresponding label. The objective is to learn a mapping $f(q,o)\rightarrow R$ that accurately predicts the reward. Specifically, this problem can be cast as a classification task where rewards serve as binary labels (e.g., correct vs. incorrect). To this end, we define the learning objective optimized via a standard binary cross-entropy loss as follows:
$\mathcal{L}(\theta)=-\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(\cdot|q)}\bigg{[}R(q,o)\log\Big{(}\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{)}+\big{(}1-R(q,o)\big{)}\log\Big{(}1-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{)}\bigg{]},$ (1)
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function that maps real-valued scores to the $[0,1]$ interval, representing the predicted probability of correctness. The score function $\psi(q,o;\pi_{\theta})$ is parameterized in terms of the policy model $\pi_{\theta}$, which is trained to estimate the quality of the generated outputs directly. The specific form and implementation of $\psi(q,o;\pi_{\theta})$ will be discussed in detail in Section 3.4.
3.3 Gradient Analysis of Objective Function
To gain a deeper understanding of how our objective guides learning, we analyze the gradient of the loss function $\mathcal{L}(\theta)$ with respect to the model parameters $\theta$. We begin by introducing the per-sample loss term:
$l(q,o;\pi_{\theta})\coloneqq R(q,o)\log\Big{(}\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{)}+\big{(}1-R(q,o)\big{)}\log\Big{(}1-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{)}$ (2)
which corresponds to the cross-entropy between the ground-truth reward $R(q,o)$ and the prediction $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$. Using this definition, the gradient of the full loss can be expressed as:
$\nabla_{\theta}\mathcal{L}(\theta)=$ $-\mathbb{E}_{q}\Big{[}\nabla_{\theta}\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\big{[}l(q,o;\pi_{\theta})\big{]}\Big{]}$ (3)
$=$ $-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\Big{[}\nabla_{\theta}\log\pi_{\theta}(o|q)\big{[}l(q,o;\pi_{\theta})\big{]}+\nabla_{\theta}l(q,o;\pi_{\theta})\Big{]}$
We now derive the inner gradient term $\nabla_{\theta}l(q,o;\pi_{\theta})$. Applying the chain rule yields:
$\nabla_{\theta}l(q,o;\pi_{\theta})=\left[R(q,o)\frac{\sigma^{\prime}\big{(}\psi(q,o;\pi_{\theta})\big{)}}{\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}}+\big{(}1-R(q,o)\big{)}\frac{-\sigma^{\prime}\big{(}\psi(q,o;\pi_{\theta})\big{)}}{1-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}}\right]\nabla_{\theta}\psi(q,o;\pi_{\theta})$ (4)
Recalling that the derivative of the sigmoid function is $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$, we simplify the gradient term as follows:
$\nabla_{\theta}l(q,o;\pi_{\theta})=\Big{[}R(q,o)-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{]}\nabla_{\theta}\psi(q,o;\pi_{\theta})$ (5)
Substituting Equation 5 into the expression for $\nabla_{\theta}\mathcal{L}(\theta)$, we obtain the final form of the gradient:
$\nabla_{\theta}\mathcal{L}(\theta)=-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\bigg{[}\underbrace{\big{[}l(q,o;\pi_{\theta})\big{]}\nabla_{\theta}\log\pi_{\theta}(o|q)}_{\text{{{Actor}}: policy improvement}}+\underbrace{\Big{(}R(q,o)-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{)}\nabla_{\theta}\psi(q,o;\pi_{\theta})}_{\text{{{Critic}}: reward estimation}}\bigg{]}$ (6)
In the above gradient expression, the first term corresponds to a standard policy gradient update, weighted by the per-sample cross-entropy loss. The second term serves as a reward estimation correction, adjusting the score function $\psi$ to better align the predicted reward $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ with the ground-truth outcome. Notably, this formulation unifies the Actor and Critic updates within a single gradient step and shared parameter space, enabling tighter coupling and more efficient learning.
Implicit Actor Critic Coupling.
As shown in Equation 6, our formulation enables the learned model $\pi_{\theta}$ to simultaneously fulfill two roles:
  • Actor: It samples outputs $o$ conditioned on the query $q$, according to the policy distribution $\pi_{\theta}(o|q)$.
  • Critic: It estimates the quality of each sampled output using the score function $\psi(q,o;\pi_{\theta})$, where the sigmoid-transformed value $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ represents the predicted probability of correctness and serves as the predicted reward.
In the Actor update, the per-sample cross-entropy loss term $l(q,o;\pi_{\theta})$ quantifies the alignment between the predicted reward $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ and the ground-truth reward $R(q,o)$. This term serves as a dynamic weight in the policy gradient, modulating the magnitude of updates based on prediction accuracy. In the Critic update, the residual term $R(q,o)-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ captures the discrepancy between the predicted and actual reward. This prediction error serves as a direct supervised learning signal, driving the score function $\psi(q,o;\pi_{\theta})$ to better predict the outcome reward. Notably, improvements in the Critic directly influence future Actor updates by refining the weighting term $l(q,o;\pi_{\theta})$.
In general, the gradient decomposition in Equation 6 splits the gradient into two components: the policy gradient update (Actor) and the supervised reward estimation updates (Critic). Crucially, both components operate within a single policy model $\pi_{\theta}$, enabling implicit coupling between Actor and Critic without the need for separate networks or alternating update schedules. This unified design facilitates efficient training while leveraging the strengths of both policy gradient and supervised learning paradigms.
3.4 Score Function Instantiation
To instantiate the score function $\psi(q,o;\pi_{\theta})$, we require a mechanism to quantify how good a sampled response $o$ is for a given query $q$. In our work, we interpret and implement $\psi(q,o;\pi_{\theta})$ as an advantage-like function, which measures the relative quality of a response conditioned on the query within a group. Unlike value-model-free methods that estimate advantages using ground-truth rewards, we define $\psi(q,o;\pi_{\theta})$ directly in terms of the policy model $\pi_{\theta}$, enabling direct policy optimization. To compute this advantage efficiently and with minimal implementation overhead, we adopt the REINFORCE Leave-One-Out (RLOO) estimator (Kool et al., 2019; Ahmadian et al., 2024). RLOO provides an unbiased estimate of the relative advantage by comparing each sampled output to others generated for the same query within a batch. Specifically, for each query $q$, we sample a set of $G$ candidate responses $o_{1},o_{2},\dots,o_{k}$}. For each output $o_{i}$, the RLOO-based advantage-like score function is defined as:
$\psi(q,o_{i};\pi_{\theta})=\hat{r}(q,o_{i};\pi_{\theta})-\frac{1}{G-1}\sum_{j\neq i}\hat{r}(q,o_{j};\pi_{\theta}),$ (7)
where $\hat{r}(q,o_{i};\pi_{\theta})$ is a reward proxy based on log-probability ratios, following prior work (Rafailov et al., 2023):
$\hat{r}(q,o_{i};\pi_{\theta})=\beta\log\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\text{ref}}(o_{i}|q)}=\beta\sum_{t=1}^{|o_{i}|}\Big{(}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})-\log\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})\Big{)},$ (8)
where $\pi_{\text{ref}}$ is a fixed reference policy used to regularize the learned policy, and $\beta$ is a scaling hyperparameter. However, in practical implementation, the reward proxy $\hat{r}(q,o_{i};\pi_{\theta})$ may grow progressively larger over time due to the fixed reference policy, leading to high variance and instability during training. To address this, we follow the strategy proposed in (Liu et al., 2025a), periodically hard-resetting the reference policy $\pi_{\text{ref}}$ to a recent snapshot of the online policy $\pi_{\theta}$, and reinitializing optimizer states to maintain stable training dynamics. With the RLOO-estimated advantage serving as the score function, the training objective becomes:
$\mathcal{L}_{\text{PACS}}(\theta)=$ $-\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot|q)}$ (9)
$\frac{1}{G}\sum_{i=1}^{G}\left[R(q,o_{i})\log\left(\sigma(\psi(q,o_{i};\pi_{\theta}))\right)+(1-R(q,o_{i}))\log\left(1-\sigma(\psi(q,o_{i};\pi_{\theta}))\right)\right].$
This formulation preserves the supervised learning structure of our original objective while incorporating relative comparisons between sampled outputs, yielding richer and more informative learning signals [^1] A rule-based reward function is employed, assigning a reward of $1$ to correct answers and 0 to incorrect ones.. The overall PACS framework is illustrated in Figure 2.
3.4.1 Loss Function Analysis.
To gain deeper insight into the learning dynamics, we analyze the loss function in Equation 9. When a sampled output $o_{i}$ is correct (i.e., $R(q,o_{i})=1$), the loss reduces to the term $\log(\sigma(\psi(q,o_{i};\pi_{\theta})))$. Minimizing the loss in this case corresponds to maximizing $\psi(q,o_{i};\pi_{\theta})$, which in turn requires increasing the estimated reward proxy $\hat{r}(q,o_{i};\pi_{\theta})$. Since $\hat{r}$ is defined as a log-probability ratio, this implies increasing $\pi_{\theta}(o_{i}|q)$, i.e., assigning higher probability to generating the correct response. In this way, the model learns to favor correct outputs. Conversely, when a sampled output $o_{i}$ is incorrect (i.e., $R(q,o_{i})=0$), the loss reduces to the term $\log(1-\sigma(\psi(q,o_{i};\pi_{\theta})))$. Minimizing the loss requires decreasing $\psi(q,o_{i};\pi_{\theta})$, which again corresponds to lowering the reward proxy $\hat{r}(q,o_{i};\pi_{\theta})$. This is achieved by reducing $\pi_{\theta}(o_{i}|q)$, discouraging the generation of incorrect responses. Thus, the model learns not only to generate correct outputs, but also to avoid producing incorrect ones.
3.4.2 Handling Data Imbalance.
To address the potential distributional imbalance in generated outputs $o_{i}$ for $q$, we adopt the class imbalance treatment methodology proposed by (King & Zeng, 2001), whereby differential weights are assigned to correct and incorrect samples respectively. Specifically, when the proportion of correct and incorrect samples in the training data exhibits imbalance, we mitigate the adverse effects of sample distribution bias on model performance by adjusting weight parameters to balance the model’s attention across different categorical samples.
MATH 500 ($\text{pass@}k$) AMC 23 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 56.66 67.64 75.75 81.91 86.87 90.60 30.20 66.16 83.88 90.42 94.85 97.10
PPO 65.48 71.94 77.51 82.14 86.02 89.20 38.57 66.07 80.19 86.81 92.09 95.53
GRPO 66.92 73.34 78.60 83.01 86.60 89.00 37.01 68.71 82.61 87.72 92.39 96.44
PACS 67.31 74.75 80.60 85.03 88.22 90.40 39.63 68.23 82.45 88.23 92.70 96.18
AIME 2024 ($\text{pass@}k$) AIME 2025 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 4.92 15.66 27.21 33.45 40.95 49.68 1.73 10.62 23.45 30.69 39.09 47.38
PPO 7.55 19.70 30.27 35.69 42.08 50.81 1.40 8.99 21.11 28.07 35.52 42.51
GRPO 8.14 19.13 26.06 31.65 37.75 42.76 2.62 7.07 14.07 20.18 28.37 36.58
PACS 7.98 21.49 30.10 35.87 42.25 48.32 2.94 11.32 22.91 30.50 39.07 47.18
Table 1: Results of Qwen2.5-3B trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best. Due to space constraints, selected results are presented here. The complete results are provided in AMC23 (Table 5), AIME 2024 (Table 7) and AIME 2025 (Table 9) in the Appendix.
4 Experiments
4.1 Experimental Setup
Datasets & Models.
This study uniformly adopts the DeepScaleR (Luo et al., 2025) dataset as the training corpus, which constitutes a high-quality mathematical problem-solving collection comprising approximately 40,000 unique mathematical problem-answer pairs. To comprehensively evaluate the performance of methods, we conduct extensive evaluations on four representative benchmarks: MATH 500, AMC23, AIME 2024, and AIME 2025. The experiments utilize Qwen2.5-3B and Qwen2.5-7B (Qwen et al., 2025) as base models, enabling us to assess PACS’s performance across various model scales.
Baselines & Hyperparameters.
To evaluate the effectiveness of our approach, we compare the performance of PACS with representative RL algorithms: PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024). The training objectives of these algorithms can be found in Appendix. We also report the performance of base models for reference. The train batch size is set to 1,024. The learning rate of actor model is set to 1e-6. For each query, $8$ responses are sampled. For PACS, $\beta$ is set to 1.0. During inference, we configure the sampling parameters with a temperature of 0.6 and top-$p$ of 0.96. For Qwen2.5-3B, we conduct training for 140 steps, while for Qwen2.5-7B, we extend the training to 300 steps. Prompt template and detailed hyperparameter configuration can be found in the Appendix.
Hardware.
All experiments are conducted on NVIDIA H200 GPUs. Training employs verl (Sheng et al., 2024), an efficient reinforcement learning library specifically designed for large language models. For inference, we utilize vllm (Kwon et al., 2023).
Evaluation Metrics.
To evaluate the model’s capability to generate diverse correct solutions while mitigating the randomness bias that may arise from single sampling, we employ the $\text{pass@}k$ metric for assessment. $\text{pass@}k$ represents the probability that at least one correct answer exists among $k$ generated candidate solutions. $\text{pass@}k$ also enables performance evaluation under different computational budgets by adjusting the value of $k$. However, directly calculating $\text{pass@}k$ with only $k$ samples per instance often results in high variance. To obtain more stable evaluation results, we generate $n$ samples for each problem (where $n\geq k$), count the number of correct samples $c$, and calculate the unbiased estimator (Chen et al., 2021):
$\text{pass@}k=\mathbb{E}_{x\sim D}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$]. (10)
Our sampling strategy adapts to the scale of dataset. For the MATH 500 dataset, given its scale of 500 problems, the sampling size is set to $n=32$ for each problem and model performance is evaluated across $k\in\{1,2,4,8,16,32\}$. Conversely, for smaller test sets (AMC23, AIME 2024, and AIME 2025), which contain fewer problems but with high-quality content, larger sampling scale is employed to obtain more stable and reliable statistical results. Specifically, we generate 512 candidate solutions for each problem ($n=512$) and conduct evaluation across a broader range of $k$ where $k\in\{1,2,4,8,16,32,64,128,256\}$.
MATH 500 ($\text{pass@}k$) AMC 23 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 61.55 73.90 82.09 87.48 90.93 92.80 35.72 70.21 85.53 91.19 95.89 99.18
PPO 76.56 80.74 84.15 86.70 88.60 90.20 51.22 72.80 81.92 85.71 88.67 90.92
GRPO 78.29 82.08 84.70 86.60 88.26 90.00 57.22 72.20 80.72 86.14 90.69 93.34
PACS 78.97 84.39 88.18 90.68 92.33 93.60 57.83 78.17 88.22 91.95 95.51 98.51
    - w/o weight 77.84 83.74 87.68 90.30 92.06 93.20 54.51 75.15 83.77 86.89 88.97 90.62
AIME 2024 ($\text{pass@}k$) AIME 2025 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 7.26 23.46 36.41 43.77 52.66 62.55 3.35 16.04 28.56 34.99 42.07 50.11
PPO 15.56 28.50 38.71 41.50 43.85 46.46 9.61 24.89 32.32 34.94 38.27 42.91
GRPO 14.10 25.37 34.78 38.50 41.66 45.42 8.36 21.63 33.26 38.12 42.33 46.45
PACS 15.80 29.75 43.90 51.21 56.54 59.78 13.33 29.65 42.88 49.45 54.48 58.22
    - w/o weight 15.72 24.57 32.58 38.77 46.16 53.42 10.44 22.86 31.03 36.59 42.39 47.67
Table 2: Results of Qwen2.5-7B trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best. Due to space constraints, selected results are presented here. The complete results are provided in AMC23 (Table 6), AIME 2024 (Table 8) and AIME 2025 (Table 10) in the Appendix.
Figure 3: Performance analysis of PACS with varying $\beta$. The 3D heatmaps show pass@k scores for different combinations of $\beta$ values (0.1, 0.5, 1, 2, 10) and $k$ values on MATH-500, AMC23, AIME-2024, and AIME-2025 datasets.
4.2 Main Results
Table 1 and Table 2 compare the overall results of PACS with different baselines. Experimental results on the Qwen2.5-3B model demonstrate that PACS exhibits significant performance advantages across multiple mathematical reasoning tasks. On the MATH 500 dataset, PACS achieves optimal performance across all $k$-value configurations, particularly attaining a pass rate of 67.31% at $\text{pass@}1$, representing an improvement of 10.65 points over the base model and surpassing both PPO and GRPO. In the AMC 23 task, PACS similarly demonstrates exceptional performance. Notably, on the more challenging AIME 2024 and AIME 2025, PACS maintains its leading position across most $k$-value settings, demonstrating the algorithm’s stability and effectiveness in complex mathematical reasoning tasks.
On the larger-scale Qwen2.5-7B model, the advantages of the PACS algorithm become even more pronounced, further validating its generalization capability across different model scales. On the MATH 500 dataset, PACS achieves superior performance across all evaluation metrics. The performance of PACS is especially remarkable on the AIME series tasks: in AIME 2024, it achieves a pass rate of 59.78% at $\text{pass@}256$, representing improvements of 13.32 and 14.36 points compared to PPO and GRPO, respectively; in AIME 2025, the $\text{pass@}256$ metric shows improvements of 15.31 and 11.77 points compared to PPO and GRPO, respectively. These results conclusively demonstrate PACS’s exceptional performance in handling high-difficulty mathematical reasoning problems and its significant advantages over reinforcement learning methods PPO and GRPO.
4.3 Ablation Study
Different $\beta$.
To validate the impact of $\beta$ on the model’s performance in Equation 8, we design and conduct ablation experiments. Models are trained with different $\beta\in\{0.1,0.5,1,2,10\}$ and their performance is evaluated in Figure 3. The results demonstrate that model achieves optimal performance when $\beta=1$, yielding higher $\text{pass@}k$.
Further analysis reveals that the AIME 2024 and AIME 2025 datasets exhibit heightened sensitivity to variations in $\beta$. This suggests that for tasks of higher complexity and difficulty, $\beta$ constitutes a critical step in achieving optimal performance. In contrast, the AMC23 and MATH 500 datasets demonstrate greater robustness across different $\beta$, with minor fluctuations in model performance. This phenomenon can be attributed to the relatively simpler nature of problems in these datasets, enabling the model to maintain stable performance across a broader range of hyperparameters.
Different Advantage Estimators.
To investigate the influence of advantage estimators on PACS’s efficacy, we extend our evaluation beyond the default RLOO to include GRPO (Shao et al., 2024) and Dr. GRPO (Liu et al., 2025b) as shown in Appendix C.4. Our findings indicate that the PACS with RLOO exhibits superior overall performance, a lead that becomes more significant at higher $\text{pass@}k$ metrics. Notably, whereas RLOO and Dr. GRPO yield comparable results on datasets such as MATH 500 and AMC23, RLOO’s superiority is markedly evident on the more demanding AIME benchmark. We posit that this performance gap stems from a subtle algorithmic divergence between RLOO and Dr. GRPO, which is magnified under more challenging conditions. By contrasting a single sample with the cohort average, RLOO’s leave-one-out mechanism yields more robust and precise signals for credit assignment.
Effect of Weight Integration.
To validate the effectiveness of the weight mechanism in PACS, we conduct ablation experiments by comparing the performance of PACS with and without the weight component. The results presented in Table 2 clearly demonstrate the positive impact of the weight mechanism. PACS consistently outperforms its weight-free variant (w/o weight). The benefits are particularly pronounced on challenging AIME datasets, where the weight component proves crucial for complex mathematical reasoning tasks, achieving substantial performance improvements especially at higher sampling rates (e.g., 6.36% point improvement on AIME 2024 at $\text{pass@}32$ of Qwen2.5-7B).
4.4 Training Dynamics
Figure 4: Training dynamics of Qwen2.5-7B on the DeepScaleR dataset across different optimization methods (PPO, GRPO, PACS, and PACS w/o weight). The plots illustrate the evolution of three key metrics during training: (a) entropy loss, (b) gradient norm, and (c) mean response length over training steps.
Figure 4 compares training dynamics of four methods applied to Qwen2.5-7B. PPO and GRPO maintain consistently low entropy loss, resulting in entropy collapse and conservative outputs. But PACS shows a non-monotonic pattern: entropy increases to about 1,700 at step 220, then stabilizes at 1,200-1,300, reflecting a superior exploration-exploitation balance. PACS w/o weight demonstrates intermediate behavior. Gradient norms reveal similar patterns. While PPO and GRPO maintain stable gradient norms throughout training, PACS peaks at 40 with sustained elevation, enabling aggressive parameter updates and broader exploration. This sustained gradient elevation in PACS, contrasting with the stable but lower gradient norms in baseline methods, demonstrates superior training efficiency. The consistently higher gradient activity indicates that PACS continues making meaningful parameter updates throughout the training process, while the stable but modest gradients in PPO and GRPO suggest more conservative optimization that may lead to premature convergence. Response lengths show PACS generating the most detailed responses. PACS achieves superior exploration capability and response quality through moderate entropy maintenance and dynamic gradient updates, providing novel optimization insights compared to other methods.
5 Conclusion
In this work, we propose PACS, a novel RLVR algorithm that achieves implicit actor-critic coupling via supervised learning. By treating outcome rewards as supervised targets and parameterizing the score function with policy log probabilities, PACS addresses the sparse reward and training instability challenges in existing RL-based RLVR methods. Extensive experiments on four mathematical benchmark datasets demonstrate that PACS significantly outperforms strong baselines including PPO and GRPO, achieving superior task performance while maintaining healthier policy entropy for sustained exploration. These findings suggest that supervised learning-based implicit actor-critic coupling represents a promising paradigm for advancing RLVR.
Appendix A Baselines
A.1 Proximal Policy Optimization Algorithms (PPO)
The objective function of PPO (Schulman et al., 2017) is formulated as follows:
$\mathcal{J}_{\text{PPO}}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta_{\text{old}}}(\cdot|q)}\frac{1}{|o|}\sum_{t=1}^{|o|}\left\{\min\left[\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}A_{t},\text{clip}\left(\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})},1-\varepsilon,1+\varepsilon\right)A_{t}\right]\right\}$
where $A$ represents the advantage function estimated using Generalized Advantage Estimation (GAE), and $\varepsilon$ denotes the clipping hyperparameter that stabilizes the training process in PPO.
A.2 Group Relative Policy Optimization (GRPO)
The objective function of GRPO (Shao et al., 2024) is defined as:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|q)}$
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}A_{i,t},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)A_{i,t}\right]-\beta D_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\right\}$
where the advantage function is computed as:
$A_{i,t}=\frac{R(q,o_{i})-\text{mean}(\{R(q,o_{1}),\dots,R(q,o_{G})\})}{\text{std}(\{R(q,o_{1}),\dots,R(q,o_{G})\})}$
The KL divergence term is estimated using:
$D_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]=\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-1.$
Following the approaches in DAPO (Yu et al., 2025) and Dr. GRPO (Liu et al., 2025b), the KL divergence term may be optionally omitted from the original GRPO objective (i.e., setting $\beta=0$), as empirical evidence suggests it may not be essential for performance.
Appendix B Hyperparameters Settings
To ensure Experimental efficiency and effectiveness, specialized optimized frameworks are used for training and inference phases. This section details the key hyperparameters used in each phase.
B.1 Template
[…]
Chat Template
<|im_start|>system
You are a helpful assistant.<|im_end|>
<|im_start|>user
{input} Please reason step by step, and put your final answer within
[…]
<|im_start|>assistant
B.2 Training Hyperparameters
During the model training phase, the verl (Sheng et al., 2024) framework is employed, which is a powerful toolkit designed for large-scale model training. The relevant hyperparameter settings during training are as follows:
Hyperparameters Configuration
Train batch size 1,024
Max prompt length 512
Max response length 4,096
Filter overlong prompts True
Mini batch size 256
Micro batch size per GPU 16 for Qwen2.5-3B, 8 for Qwen2.5-7B
Learning rate $1\times 10^{-6}$ for actor, $1\times 10^{-5}$ for critic in PPO
Rollout number 8
Use KL loss False
Total training steps 140 for Qwen2.5-3B, 300 for Qwen2.5-7B
Table 3: Training hyperparameters.
B.3 Inference Hyperparameters
During the model inference and evaluation phase, the high-performance inference serving framework vllm (Kwon et al., 2023) is employed. Through vllm, high-throughout and low-latency text generation is achieved. Consistent inference configurations are adopted across all experiments to ensure fairness and comparability of evaluation results.
Hyperparameters Configuration
Enable prefix caching True
GPU memory utilization 0.9
Max tokens 4,096
Temperature 0.6
Top-pp 0.95
Tensor parallel size 4
Table 4: Inference hyperparameters.
Appendix C Detailed Experimental Results
This section presents the comprehensive test results of each model on AMC 23, AIME 2024, and AIME 2025, along with corresponding analyses.
C.1 AMC23
The experimental results clearly demonstrate the superior performance of the PACS algorithm on the AMC23. As shown in Table 5 and Table 6, PACS achieves optimal or near-optimal performance across the majority of $\text{pass@}k$ metrics. The advantages of PACS are particularly pronounced at lower $k$ values; for instance, on the Qwen2.5-3B and Qwen2.5-7B models, its pass@1 scores reach 39.63 and 57.83 respectively, surpassing all baselines. This indicates that the proposed algorithm can more effectively enhance the model’s capability to solve problems in the first attempt or within a few trials, which holds significant value in practical applications. As the model scale expands from 3B to 7B, the PACS algorithm demonstrates excellent scalability, with its performance advantages becoming increasingly pronounced. On the Qwen2.5-7B model, PACS not only achieves leading performance in pass@1 but also comprehensively outperforms other baselines across the broad range from $k=1$ to $k=64$, showing substantial performance improvements. For example, at pass@8, PACS (78.17) outperforms the second-best GRPO (72.20) by nearly 6 points.
AMC23 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 30.20 42.56 55.10 66.16 75.62 83.88 90.42 94.85 97.10
PPO 38.57 48.65 57.97 66.07 73.25 80.19 86.81 92.09 95.53
GRPO 37.01 46.85 58.14 68.71 76.69 82.61 87.72 92.39 96.44
PACS 39.63 50.56 60.30 68.23 75.97 82.45 88.23 92.70 96.18
Table 5: Results of Qwen2.5-3B on AMC23 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
AMC23 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 35.72 49.47 61.05 70.21 78.35 85.53 91.19 95.89 99.18
PPO 51.22 60.01 67.22 72.80 77.52 81.92 85.71 88.67 90.92
GRPO 57.22 63.36 68.21 72.20 76.06 80.72 86.14 90.69 93.34
PACS 57.83 65.57 71.94 78.17 83.83 88.22 91.95 95.51 98.51
    - w/o weight 54.51 62.95 69.65 75.15 79.77 83.77 86.89 88.97 90.62
Table 6: Results of Qwen2.5-7B on AMC23 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
C.2 AIME 2024
On the more challenging AIME 2024 dataset, PACS again demonstrates its effectiveness. On the Qwen2.5-3B model scale, different algorithms show very similar performance with intense competition, yet PACS still achieves leading results in critical intervals. Throughout the entire $\text{pass@}k$ curve, PACS consistently remains in the top tier. On the Qwen2.5-7B model, PACS achieves comprehensive leadership from $k=1$ to $k=128$, obtaining optimal results at nearly all evaluation points. Similar to the AMC experiments, the base model again achieves the highest score at $k=256$, which further confirms the inference that unaligned base models retain higher answer diversity and have the potential to “enumerate” correct answers through massive sampling, though this approach is highly inefficient for practical applications.
AIME 2024 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 4.92 7.90 11.43 15.66 21.11 27.21 33.45 40.95 49.68
PPO 7.55 10.83 14.96 19.70 24.94 30.27 35.69 42.08 50.81
GRPO 8.14 11.45 14.79 18.13 21.74 26.06 31.65 37.75 42.76
PACS 7.98 12.28 16.98 21.49 25.58 30.10 35.87 42.25 48.32
Table 7: Results of Qwen2.5-3B on AIME 2024 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
AIME 2024 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 7.26 11.66 17.17 23.46 29.94 36.41 43.77 52.66 62.55
PPO 15.56 19.45 23.51 28.50 34.16 38.71 41.50 43.85 46.46
GRPO 14.10 16.92 20.71 25.37 30.34 34.78 38.50 41.66 45.42
PACS 15.80 19.93 24.39 29.75 36.29 43.90 51.21 56.54 59.78
    - w/o weight 15.72 19.60 22.24 24.57 27.88 32.58 38.77 46.16 53.42
Table 8: Results of Qwen2.5-7B on AIME 2024 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
C.3 AIME 2025
To comprehensively evaluate the performance stability and generalization capability of PACS, we conduct assessments on the latest AIME 2025 benchmark. This dataset represents the current highest difficulty standard for mathematical reasoning tasks, enabling more rigorous examination of algorithm performance under extremely challenging scenarios. On 3B-scale models, PACS demonstrates impressive comprehensive advantages, achieving optimal performance across nearly all evaluation metrics. PACS obtains superior results throughout the entire evaluation spectrum from pass@1 to pass@256.
On 7B models, PACS achieves complete dominance, from pass@1 of 13.33 to pass@256 of 58.22, with every metric substantially outperforming all baseline methods. It is worth emphasizing that the performance gap between PACS and the second-ranked method is highly significant across multiple key metrics. For instance, at pass@16, PACS (36.00) surpasses GRPO (27.57) by over 8 percentage points; at pass@128, the gap reached 12 percentage points (54.48 vs 42.33).
AIME 2025 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 1.73 3.32 6.14 10.62 16.67 23.45 30.69 39.09 47.38
PPO 1.40 2.71 5.07 8.99 14.56 21.11 28.07 35.52 43.51
GRPO 2.62 3.92 5.29 7.07 9.87 14.07 20.18 28.37 36.58
PACS 2.94 4.92 7.63 11.32 16.47 22.91 30.50 39.07 47.18
Table 9: Results of Qwen2.5-3B on AIME 2025 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
AIME 2025 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 3.35 6.12 10.40 16.04 22.32 28.56 34.99 42.07 50.11
PPO 9.61 14.42 19.69 24.89 29.21 32.32 34.94 38.27 42.91
GRPO 8.36 11.82 16.34 21.63 27.57 33.26 38.12 42.33 46.45
PACS 13.33 18.19 23.73 29.65 36.00 42.88 49.45 54.48 58.22
    - w/o weight 10.44 15.12 19.36 22.86 26.48 31.03 36.59 42.39 47.67
Table 10: Results of Qwen2.5-7B on AIME 2025 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
Synthesizing results across the four benchmarks—MATH 500, AMC23, AIME 2024, and AIME 2025—PACS exhibits exceptional cross-task generalization capability and performance consistency. Whether on the relatively straightforward MATH 500 and AMC23 or the extremely challenging AIME 2025, PACS consistently outperforms existing mainstream methods. This conclusively demonstrates that our proposed algorithm possesses robust adaptability and resilience, enabling outstanding performance across mathematical reasoning tasks of varying difficulty levels.
C.4 Ablation Study
To conduct a comparative analysis of how different advantage estimators affect PACS performance, we benchmark our default RLOO method against two alternatives: GRPO (Shao et al., 2024) and Dr. GRPO (Liu et al., 2025b). While RLOO utilizes a leave-one-out mechanism, the advantage functions for GRPO and Dr. GRPO are defined as follows. Dr. GRPO introduces simple yet significant modifications to address the biases in GRPO by removing the std normalization terms.
$\psi_{\text{GRPO}}(q,o_{i};\pi_{\theta})=\frac{\hat{r}(q,o_{i};\pi_{\theta})-\text{mean}(\{\hat{r}(q,o;\pi_{\theta})\})}{\text{std}(\{\hat{r}(q,o;\pi_{\theta})\})},$ (11)
$\psi_{\text{Dr. GRPO}}(q,o_{i};\pi_{\theta})=\hat{r}(q,o_{i};\pi_{\theta})-\text{mean}(\{\hat{r}(q,o;\pi_{\theta})\}),$ (12)
MATH 500 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$
PACS 78.97 84.39 88.18 90.68 92.33 93.60
    - GRPO 74.71 80.19 84.31 87.42 89.64 91.20
    - Dr. GRPO 79.68 84.44 87.90 90.52 92.30 93.40
Table 11: Results of different advantage estimators of Qwen2.5-7B on MATH 500. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
AMC23 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
PACS 57.83 65.57 71.94 78.17 83.83 88.22 91.95 95.51 98.51
    - GRPO 48.66 57.43 63.96 69.68 75.71 81.63 86.94 90.95 94.06
    - Dr. GRPO 58.29 66.54 72.83 78.27 83.28 87.45 90.12 91.65 93.13
Table 12: Results of different advantage estimators of Qwen2.5-7B on AMC23. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
On the MATH 500 and AMC23 benchmarks (see Tables 11 and  12), PACS (employing RLOO) demonstrates highly comparable performance with the Dr. GRPO variant. This result aligns with expectations, as RLOO and Dr. GRPO exhibit only subtle algorithmic differences, both enhancing the foundational GRPO through intra-batch relative comparisons. On these moderately challenging tasks, such minor algorithmic distinctions do not translate into substantial performance disparities. Specifically, Dr. GRPO shows marginal superiority at lower $k$ (kk=1, 2), while PACS exhibits superior performance at higher $k$ values; however, the overall performance of both methods remains remarkably similar.
AIME 2024 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
PACS 15.80 19.93 24.39 29.75 36.29 43.90 51.21 56.54 59.78
    - GRPO 12.75 16.24 19.00 21.59 24.88 28.91 33.22 37.14 39.95
    - Dr. GRPO 15.59 19.67 23.99 29.04 34.78 40.82 46.76 53.08 59.68
Table 13: Results of different advantage estimators of Qwen2.5-7B on AIME 2024. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
AIME 2025 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
PACS 13.33 18.19 23.73 29.65 36.00 42.88 49.45 54.48 58.22
    - GRPO 8.39 13.01 17.69 21.78 25.77 29.86 34.34 39.26 43.73
    - Dr. GRPO 13.93 17.15 20.23 24.05 28.68 34.55 41.07 46.12 49.56
Table 14: Results of different advantage estimators of Qwen2.5-7B on AIME 2025. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
The superiority of RLOO is most pronounced on the AIME 24 and AIME 25 (Tables 13 and 14). Here, the nuanced algorithmic difference between RLOO and Dr. GRPO is amplified into a clear performance divergence, with our PACS algorithm demonstrating unequivocal superiority across all values of $k$. We posit that this performance gap is rooted in the distinctive ”leave-one-out” credit assignment of RLOO. High-difficulty tasks with sparse rewards and high solution variance demand a robust signal for policy updates. RLOO’s mechanism—contrasting each sample against the cohort average—provides exactly this stability. The practical consequence is substantial: on the AIME 2025 benchmark, for instance, PACS (58.22%) establishes a lead of almost 9 percentage points over Dr. GRPO (49.56%) at pass@256.

### v2

@@
 Recent advances in Reinforcement Learning with Verifiable Rewards (RLVR) have empowered large language models (LLMs) to tackle challenging reasoning tasks such as mathematics and programming.
-RLVR leverages verifiable outcome rewards to guide policy optimization, enabling LLMs to progressively improve output quality in a grounded and reliable manner.
-Despite its promise, the RLVR paradigm poses significant challenges, as existing methods often suffer from sparse reward signals and unstable policy gradient updates, particularly in RL-based approaches.
+Despite its promise, the RLVR paradigm poses significant challenges, as existing methods often suffer from sparse reward signals and unstable policy gradient updates, inherent to RL-based approaches.
 To address the challenges, we propose PACS, a novel RLVR framework that achieves imPlicit Actor Critic coupling via a Supervised learning framework.
 By treating the outcome reward as a predictable label, we reformulate the RLVR problem into a supervised learning task over a score function parameterized by the policy model and optimized using cross-entropy loss.
-A detailed gradient analysis shows that this supervised formulation inherently recovers the classical policy gradient update while implicitly coupling actor and critic roles, yielding more stable and efficient training.
-Benchmarking on challenging mathematical reasoning tasks, PACS outperforms strong RLVR baselines, such as PPO and GRPO, achieving superior reasoning performance.
-For instance, PACS achieves 59.78% at pass@256 on AIME 2025, representing improvements of 13.32 and 14.36 points over PPO and GRPO.
-This simple yet powerful framework offers a promising avenue for LLMs post-training with verifiable rewards.
+A detailed gradient analysis shows that this supervised formulation inherently recovers the classical policy gradient update while providing more stable and efficient training.
+Extensive experiments demonstrate that PACS significantly outperforms strong open-source models and RLVR baselines, yielding substantial average gains of +8.26% (4B) and +9.57% (8B) over base models offering a promising avenue for LLMs post-training with verifiable rewards.
 Our code and data are available as open source at https://github.com/ritzz-ai/PACS.
 1 Introduction
-Large Language Models (LLMs) have recently achieved remarkable success in complex reasoning tasks, driven by advancements, such as OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025).
-In particular, scaling test-time compute enables longer Chains of Thought (CoT), facilitating more sophisticated and structured reasoning.
-This progress has led to significant improvements in challenging domains, such as mathematics (Shao et al., 2024; Liu et al., 2025a; Zuo et al., 2025) and programming (Lambert et al., 2024; Cui et al., 2025; Zhao et al., 2025).
-A crucial underlying ingredient behind these achievements is the emergence of Reinforcement Learning with Verifiable Rewards (RLVR), which empowers LLMs to progressively enhance their capabilities by reinforcing their self-generated outputs with verifiable outcome rewards.
-Existing RLVR methods fall into two primary categories: value-model-based approaches, such as PPO (Schulman et al., 2017) and VAPO (Yue et al., 2025), which employ large value models to estimate advantages for policy updates; and value-model-free approaches, such as GRPO (Shao et al., 2024; Guo et al., 2025) and DAPO (Yu et al., 2025), which bypass explicit value modeling by leveraging Monte Carlo estimation of group relative rewards.
-While both classes have demonstrated remarkable success, they face fundamental challenges stemming from the sparse nature of outcome rewards in RLVR settings, where only a single reward signal is provided after the entire response is generated.
-This sparsity severely hinders credit assignment and gradient propagation in Reinforcement Learning (RL)-based policy learning, which relies on attributing feedback to individual tokens or actions within the output.
-Value-model-based methods address this challenge by learning explicit value models, but at the cost of increased model complexity and computational overhead.
-In contrast, value-model-free methods avoid this complexity by forgoing explicit value prediction, but suffer from high variance in their Monte Carlo estimates, often leading to issues such as advantage collapse, training instability, and degraded performance.
-Figure 1: Comparison between RLVR and the supervised learning reformulation, where the query and output are input, and the outcome reward is treated as a predictable label.
+Recent advancements like OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025) have revolutionized complex reasoning by scaling test-time compute to enable longer Chains of Thought (CoT).
+This progress significantly improves performance in mathematics (Shao et al., 2024; Liu et al., 2025a; Zuo et al., 2025) and programming (Lambert et al., 2024; Cui et al., 2025; Zhao et al., 2025).
+These achievements are largely driven by Reinforcement Learning with Verifiable Rewards (RLVR), which empowers LLMs to self-improve using verifiable outcome rewards.
+Existing RLVR methods generally fall into two categories: value-model-based (e.g., PPO (Schulman et al., 2017), VAPO (Yue et al., 2025)) and value-model-free (e.g., GRPO (Shao et al., 2024; Guo et al., 2025), DAPO (Yu et al., 2025)).
+While both classes have demonstrated remarkable success, they face fundamental challenges stemming from the sparse nature of outcome rewards in RLVR settings, where only a single reward signal is provided after the entire response is generated..
+Value-model-based approaches mitigate sparsity via explicit value modeling but incur high computational overhead.
+Conversely, value-model-free methods leverage Monte Carlo estimation to avoid this cost but suffer from high variance, often leading to advantage collapse and training instability.
 This challenge, rooted in the inherently sparse nature of RLVR feedback, highlights fundamental limitations of the existing RL-based paradigm.
 It motivates the development of alternative policy optimization strategies that can leverage direct supervision to enable more stable and efficient learning in RLVR settings.
-To this end, we propose a novel framework that recasts the RLVR problem as a supervised learning task (see Figure 1).
-Instead of relying on sparse outcome rewards to optimize the policy via RL, we directly interpret the outcome reward as a supervision signal and train a score function, parameterized by the policy itself, to predict this reward via a cross-entropy loss.
-Through a detailed gradient analysis, we show that this supervised learning formulation not only recovers the standard policy gradient update but also implicitly couples the actor (policy improvement) and critic (reward estimation) components within a single update.
-Unlike conventional actor critic methods that maintain a separate value function estimator, our approach adopts a shared parameterization, allowing both roles to be updated jointly via a unified learning signal.
-This tight coupling eliminates the temporal mismatch between reward estimation and policy updates, thereby improving training efficiency.
-Moreover, in contrast to value-model-free methods, which rely solely on high-variance Monte Carlo estimates of relative advantages for policy updates, our method leverages prediction-error-based learning signals, providing more stable and reliable guidance for optimization.
-We refer to our method as PACS (imPlicit Actor Critic coupling via a Supervised learning framework).
-It offers a principled and efficient training paradigm that unifies policy learning and reward estimation within a coherent supervised learning framework.
-To assess the effectiveness of our method, we conduct experiments on four challenging mathematical reasoning tasks and compare it against the state-of-the-art RLVR methods, including PPO and GRPO.
-The results demonstrate that PACS significantly outperforms baselines.
-For instance, PACS achieves a pass@256 rate of 59.78% on AIME 2025, representing improvements of 13.32 and 14.36 points over PPO and GRPO.
-Ablation studies confirm the critical importance of the weighting mechanism, while training dynamics analysis reveals that PACS achieves superior exploration-exploitation balance.
+To address the limitations of the RL-based paradigm rooted in sparse feedback, we propose PACS (imPlicit Actor Critic coupling via a Supervised learning framework).
+We recast RLVR as a supervised learning task (as illustrated in fig.
+1) by treating outcome rewards as labels to train a policy-parameterized score function via cross-entropy loss.
+Gradient analysis reveals that this formulation inherently recovers the standard policy gradient while implicitly coupling the actor and critic through shared parameterization.
+This design eliminates the temporal mismatch introduced by separate value estimators and mitigates the high-variance issues associated with Monte Carlo estimates by providing stable prediction-error signals for optimization.
+Moreover, it reduces compute overhead compared to value-model-based training, and attenuates extreme updates that cause entropy collapse via bounded supervised residual.
+Overall, PACS offers a principled and efficient training paradigm that unifies policy learning and reward estimation within a coherent supervised learning framework.
+Figure 1: Comparison between RLVR and the supervised learning reformulation, where the query and output are input, and the outcome reward is treated as a predictable label.
+Experimental results on five benchmarks demonstrate PACS’s superiority.
+It outperforms strong open-source models like LIMO and Eurus-2-7B with a 63.32% $\text{pass@}1$ accuracy.
+Notably, PACS 4B model (45.63% on AIME 2025) surpasses the significantly larger Qwen3-14B (36.22%).
+Against RLVR baselines, PACS establishes a decisive lead, particularly on complex tasks: on BeyondAIME, PACS-8B achieves 64.02% ($\text{pass@}64$), exceeding GRPO (57.64%) and PPO (53.80%) by approximately 6.4 and 10.2 points, respectively.
+Furthermore, PACS mitigates entropy collapse, achieving significantly higher solution diversity than GRPO (0.0392 vs.
+0.0249).
 The main contributions are summarized as follows:
-  • We propose PACS, a novel RLVR approach that reframes the problem as supervised learning: motivated by the sparsity of outcome rewards, we treat them as target labels and train the policy via cross-entropy loss.
-  • We demonstrate that the proposed loss inherently captures policy gradient updates and implicitly unifies actor (policy improvement) and critic (reward estimation) through shared parameterization, enabling efficient and stable policy optimization.
-  • Extensive experiments on challenging mathematical reasoning benchmarks show that PACS outperforms state-of-the-art RLVR methods, achieving superior reasoning performance, maintaining healthy policy entropy for exploration, and exhibiting improved training efficiency.
+  • We propose PACS, a novel RLVR framework that recasts sparse outcome rewards as supervised signals, effectively training the policy via cross-entropy loss.
+  • We theoretically demonstrate that the proposed loss inherently captures policy gradient updates and implicitly unifies actor and critic through shared parameterization, enabling efficient and stable policy optimization.
+  • Extensive experiments on challenging mathematical reasoning benchmarks show that PACS consistently outperforms state-of-the-art RLVR methods, achieving superior reasoning performance.
+  • Generalization and diversity analyses demonstrate that PACS possesses strong transfer capabilities on out-of-domain tasks and effectively mitigates entropy collapse, fostering diverse reasoning paths that are instrumental to the model’s superior performance.
 2 Related Work
 Reasoning Models.
-The emergence of reasoning models, referring to LLMs explicitly trained or optimized for multi-step chain-of-thought inference, has been catalyzed by the release of OpenAI-o1 (Jaech et al., 2024), which achieves strong performance on challenging reasoning tasks.
-DeepSeek-R1 (Guo et al., 2025) further highlights the power of verifiable reward signals by using GRPO (Shao et al., 2024) with rule-based outcome rewards to guide effective policy optimization.
-Building on this foundation, QWQ (Qwen-Team, 2025) introduces a two-stage RL framework: a “cold-start” phase to bootstrap reasoning ability, followed by a brief phase of general RL finetuning to expand the model’s problem-solving proficiency.
-Most recently, Gemini 2.5 Pro (Comanici et al., 2025) integrates human feedback, retrieval-augmented generation, and multimodal perception, enabling reasoning grounded in diverse information sources beyond text.
-Together, these advances illustrate the evolution of RL-enhanced reasoning models from single-domain systems to general-purpose, multimodal agents.
+The emergence of reasoning models was catalyzed by OpenAI-o1 (Jaech et al., 2024), which demonstrated that reasoning performance can scale significantly via test-time compute.
+This paradigm was further advanced by DeepSeek-R1 (Guo et al., 2025), which utilized GRPO (Shao et al., 2024) to show that rule-based rewards can elicit self-correction and reflection without dense human supervision.
+This paradigm has been extended through the Qwen family (Team, 2025b; Qwen et al., 2025; Yang et al., 2025) and the Kimi series (Team et al., 2025c; b), which have brought high-performance reasoning to the ecosystem.
+Most recently, Gemini family (Comanici et al., 2025) and the Llama-Nemotron series (Grattafiori et al., 2024; Bercovich et al., 2025) have integrated multimodal perception and instruction-following into the reasoning loop, establishing a new state-of-the-art for general-purpose agents.
 Reinforcement Learning with Verifiable Reward (RLVR).
-RLVR has become a widely adopted strategy for enhancing the reasoning capabilities of LLMs in domains with clearly verifiable correctness, such as mathematics and programming  (Guo et al., 2025; Team et al., 2025; Ma et al., 2025; Chu et al., 2025).
+RLVR has become a widely adopted strategy for enhancing the reasoning capabilities of LLMs in domains with clearly verifiable correctness, such as mathematics and programming  (Guo et al., 2025; Ma et al., 2025; Chu et al., 2025; Yan et al., 2025; Zheng et al., 2025; Hao et al., 2025).
 Recent work has explored a wide range of RLVR techniques, which can be broadly categorized as value-model-based or value-model-free.
 Value-model-based methods, such as PPO (Schulman et al., 2017), VinePPO (Kazemnejad et al., 2024), and VAPO (Yue et al., 2025), explicitly learn a value function to estimate the expected cumulative reward, providing stable training signals at the cost of additional computational overhead.
-On the other hand, value-model-free methods, such as GRPO (Shao et al., 2024), REINFORCE++ (Hu et al., 2025), and DAPO (Yu et al., 2025), avoid explicit value modeling by relying on Monte Carlo-based advantage estimation.
-While these approaches reduce modeling complexity, they often suffer from high gradient variance, which undermines training stability and overall performance.
+On the other hand, value-model-free methods, such as GRPO (Shao et al., 2024), REINFORCE++ (Hu et al., 2025), and DAPO (Yu et al., 2025), avoid explicit value modeling by relying on Monte Carlo advantage estimation.
+While these approaches reduce modeling complexity, they often suffer from high gradient variance, which undermines training stability and performance.
 3 Methodology
-In this section, we first establish the preliminaries for RLVR, and then introduce PACS, a novel RLVR framework that achieves imPlicit Actor–Critic coupling through a Supervised learning framework.
-We present the overall framework, followed by a detailed gradient analysis that illustrates the effectiveness of our approach, and conclude with practical design considerations to instantiate the implementation.
+In this section, we first establish the preliminaries for RLVR before introducing PACS, a novel framework achieving imPlicit Actor–Critic coupling via Supervised learning.
+We then detail the overall framework, validate its effectiveness through gradient analysis, and conclude with practical implementation designs.
 3.1 Preliminaries
@@
 Various RLVR approaches (Schulman et al., 2017; Kool et al., 2019; Ahmadian et al., 2024; Shao et al., 2024; Yu et al., 2025; Yue et al., 2025) have been proposed to optimize this objective function, each seeking to effectively learn the policy $\pi_{\theta}$ under the supervision of outcome rewards.
-3.2 Recasting RLVR within a Supervised Learning Framework
-In the RLVR framework, LLMs generate responses to input queries, and verifiable outcome rewards that indicate the correctness of the final outcomes are provided at the end of responses.
-A key characteristic of this setting is that the reward is delivered only once, after the entire response is generated, reflecting the overall quality of the output.
-Inspired by this observation, we propose an alternative to conventional RL-based optimization: rather than policy learning through RL from sparse reward signals, we recast the problem as a supervised learning task in which the model is trained to directly predict the outcome reward.
 Figure 2: An illustration of the PACS framework.
@@
 (3) Cross-Entropy Loss, which converts the RLVR problem into a supervised learning task, optimizing a scoring function parameterized by the policy with a cross-entropy loss.
+3.2 Recasting RLVR via Supervised Learning
+In the RLVR framework, LLMs receive verifiable outcome rewards only after generating a complete response.
+Inspired by this observation, we propose an alternative to conventional RL-based optimization: rather than policy learning through RL from sparse reward signals, we recast the problem as a supervised learning task in which the model is trained to directly predict the outcome reward.
 Concretely, instead of using the reward $R(q,o)\in\{0,1\}$ associated with a query-output pair $(q,o)$ to guide policy optimization via RL, we consider $(q,o)$ as input data and treat $R(q,o)$ as the corresponding label.
@@
 incorrect).
-To this end, we define the learning objective optimized via a standard binary cross-entropy loss as follows:
+To this end, we define the learning objective via a standard binary cross-entropy loss as follows:
 $\mathcal{L}(\theta)=-\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(\cdot|q)}\bigg[R(q,o)\log\Big(\sigma\big(\psi(q,o;\pi_{\theta})\big)\Big)+\big(1-R(q,o)\big)\log\Big(1-\sigma\big(\psi(q,o;\pi_{\theta})\big)\Big)\bigg],$ (1)
@@
 $=$ $-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\Big[\nabla_{\theta}\log\pi_{\theta}(o|q)\big[l(q,o;\pi_{\theta})\big]+\nabla_{\theta}l(q,o;\pi_{\theta})\Big]$
-We now derive the inner gradient term $\nabla_{\theta}l(q,o;\pi_{\theta})$. Applying the chain rule yields:
-$\nabla_{\theta}l(q,o;\pi_{\theta})=\left[R(q,o)\frac{\sigma^{\prime}\big{(}\psi(q,o;\pi_{\theta})\big{)}}{\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}}+\big{(}1-R(q,o)\big{)}\frac{-\sigma^{\prime}\big{(}\psi(q,o;\pi_{\theta})\big{)}}{1-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}}\right]\nabla_{\theta}\psi(q,o;\pi_{\theta})$ (4)
-Recalling that the derivative of the sigmoid function is $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$, we simplify the gradient term as follows:
-$\nabla_{\theta}l(q,o;\pi_{\theta})=\Big{[}R(q,o)-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{]}\nabla_{\theta}\psi(q,o;\pi_{\theta})$ (5)
-Substituting Equation 5 into the expression for $\nabla_{\theta}\mathcal{L}(\theta)$, we obtain the final form of the gradient:
-$\nabla_{\theta}\mathcal{L}(\theta)=-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\bigg{[}\underbrace{\big{[}l(q,o;\pi_{\theta})\big{]}\nabla_{\theta}\log\pi_{\theta}(o|q)}_{\text{{{Actor}}: policy improvement}}+\underbrace{\Big{(}R(q,o)-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}\Big{)}\nabla_{\theta}\psi(q,o;\pi_{\theta})}_{\text{{{Critic}}: reward estimation}}\bigg{]}$ (6)
+We now derive the inner gradient term $\nabla_{\theta}\ell(q,o;\pi_{\theta})$.
+Applying the chain rule and recalling that the derivative of the sigmoid function is $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$, we simplify the gradient term as follows:
+$\nabla_{\theta}\ell(q,o;\pi_{\theta})=\Big[R(q,o)-\sigma\big(\psi(q,o;\pi_{\theta})\big)\Big]\nabla_{\theta}\psi(q,o;\pi_{\theta}).$ (4)
+Substituting eq. 4 into the expression for $\nabla_{\theta}\mathcal{L}(\theta)$, we obtain the final form of the gradient:
+$\begin{aligned} \nabla_{\theta}\mathcal{L}(\theta)=&-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\\ &\bigg[\underbrace{\big[\ell(q,o;\pi_{\theta})\big]\nabla_{\theta}\log\pi_{\theta}(o|q)}_{\text{{{Actor}}: policy improvement}}+\underbrace{\Big(R(q,o)-\sigma\big(\psi(q,o;\pi_{\theta})\big)\Big)\nabla_{\theta}\psi(q,o;\pi_{\theta})}_{\text{{{Critic}}: reward estimation}}\bigg].\end{aligned}$ (5)
 In the above gradient expression, the first term corresponds to a standard policy gradient update, weighted by the per-sample cross-entropy loss.
-The second term serves as a reward estimation correction, adjusting the score function $\psi$ to better align the predicted reward $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ with the ground-truth outcome.
-Notably, this formulation unifies the Actor and Critic updates within a single gradient step and shared parameter space, enabling tighter coupling and more efficient learning.
+The second term serves as a reward estimation correction, adjusting the score function $\psi$ to align the predicted reward $\sigma\big(\psi(q,o;\pi_{\theta})\big)$ with the ground-truth outcome.
+Notably, this formulation unifies the Actor and Critic updates within a single gradient step and shared parameter space.
 Implicit Actor Critic Coupling.
-As shown in Equation 6, our formulation enables the learned model $\pi_{\theta}$ to simultaneously fulfill two roles:
+As shown in eq. 5, our formulation enables the learned model $\pi_{\theta}$ to simultaneously fulfill two roles:
   • Actor: It samples outputs $o$ conditioned on the query $q$, according to the policy distribution $\pi_{\theta}(o|q)$.
-  • Critic: It estimates the quality of each sampled output using the score function $\psi(q,o;\pi_{\theta})$, where the sigmoid-transformed value $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ represents the predicted probability of correctness and serves as the predicted reward.
-In the Actor update, the per-sample cross-entropy loss term $l(q,o;\pi_{\theta})$ quantifies the alignment between the predicted reward $\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ and the ground-truth reward $R(q,o)$.
-This term serves as a dynamic weight in the policy gradient, modulating the magnitude of updates based on prediction accuracy.
-In the Critic update, the residual term $R(q,o)-\sigma\big{(}\psi(q,o;\pi_{\theta})\big{)}$ captures the discrepancy between the predicted and actual reward.
-This prediction error serves as a direct supervised learning signal, driving the score function $\psi(q,o;\pi_{\theta})$ to better predict the outcome reward.
-Notably, improvements in the Critic directly influence future Actor updates by refining the weighting term $l(q,o;\pi_{\theta})$.
-In general, the gradient decomposition in Equation 6 splits the gradient into two components: the policy gradient update (Actor) and the supervised reward estimation updates (Critic).
-Crucially, both components operate within a single policy model $\pi_{\theta}$, enabling implicit coupling between Actor and Critic without the need for separate networks or alternating update schedules.
-This unified design facilitates efficient training while leveraging the strengths of both policy gradient and supervised learning paradigms.
+  • Critic: It estimates output quality using the score function $\psi(q,o;\pi_{\theta})$, where the sigmoid-transformed value $\sigma\big(\psi(q,o;\pi_{\theta})\big)$ represents the predicted probability of correctness and serves as the predicted reward.
+The Actor update is modulated by the per-sample cross-entropy loss $\ell(q,o;\pi_{\theta})$, which weights the policy gradient according to the alignment between predicted and ground-truth rewards.
+The Critic update is driven by the residual $R(q,o)-\sigma\big(\psi(q,o;\pi_{\theta})\big)$, providing a supervised learning signal that improves reward prediction.
+Improving the Actor changes the score $\psi$ and thus yields a tighter supervised target for the Critic; in turn, the Critic improvements feed back to directly affect future Actor updates through the shared weighting term.
+This unified design enables implicit Actor Critic coupling without separate networks or alternating update schedules and facilitates efficient training while leveraging the strengths of both policy gradient and supervised learning paradigms.
 3.4 Score Function Instantiation
-To instantiate the score function $\psi(q,o;\pi_{\theta})$, we require a mechanism to quantify how good a sampled response $o$ is for a given query $q$.
-In our work, we interpret and implement $\psi(q,o;\pi_{\theta})$ as an advantage-like function, which measures the relative quality of a response conditioned on the query within a group.
-Unlike value-model-free methods that estimate advantages using ground-truth rewards, we define $\psi(q,o;\pi_{\theta})$ directly in terms of the policy model $\pi_{\theta}$, enabling direct policy optimization.
-To compute this advantage efficiently and with minimal implementation overhead, we adopt the REINFORCE Leave-One-Out (RLOO) estimator (Kool et al., 2019; Ahmadian et al., 2024).
-RLOO provides an unbiased estimate of the relative advantage by comparing each sampled output to others generated for the same query within a batch.
-Specifically, for each query $q$, we sample a set of $G$ candidate responses $o_{1},o_{2},\dots,o_{k}$}.
-For each output $o_{i}$, the RLOO-based advantage-like score function is defined as:
+To instantiate the score function $\psi(q,o;\pi_{\theta})$, we require a mechanism to quantify the quality of a sampled response $o$ for a given query $q$.
+We interpret $\psi(q,o;\pi_{\theta})$ as an advantage-like function that measures the relative quality of a response within a group of samples.
+Such a formulation is particularly appropriate for our setting because it centers the score distribution with respect to the set of samples while preserving its full range.
+As a result, it is well suited with a sigmoid mapping, yielding outputs confined to the $[0,1]$ interval.
+Unlike value-model-free methods that estimate advantages using ground-truth rewards, we define $\psi(q,o;\pi_{\theta})$ in terms of the policy model $\pi_{\theta}$, enabling direct policy optimization.
+To compute this advantage efficiently, we adopt the REINFORCE Leave-One-Out (RLOO) estimator (Kool et al., 2019; Ahmadian et al., 2024).
+RLOO provides an unbiased estimate of the relative advantage by comparing each sampled output to others generated for the same query.
+For each query $q$, given a set of $G$ candidate responses $\{o_{1},o_{2},\dots,o_{k}\}$, the RLOO-based advantage-like score function for each output $o_{i}$ is defined as:
 $\psi(q,o_{i};\pi_{\theta})=\hat{r}(q,o_{i};\pi_{\theta})-\frac{1}{G-1}\sum_{j\neq i}\hat{r}(q,o_{j};\pi_{\theta}),$ (6)
-where $\hat{r}(q,o_{i};\pi_{\theta})$ is a reward proxy based on log-probability ratios, following prior work (Rafailov et al., 2023):
+following prior work (Rafailov et al., 2023; Cui et al., 2025), where $\hat{r}(q,o_{i};\pi_{\theta})$ is a reward proxy based on log-probability ratios:
 $\hat{r}(q,o_{i};\pi_{\theta})=\beta\log\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\text{ref}}(o_{i}|q)}=\beta\sum_{t=1}^{|o_{i}|}\big(\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})-\log\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})\big),$ (7)
@@
 $\frac{1}{G}\sum_{i=1}^{G}\left[R(q,o_{i})\log\left(\sigma(\psi(q,o_{i};\pi_{\theta}))\right)+(1-R(q,o_{i}))\log\left(1-\sigma(\psi(q,o_{i};\pi_{\theta}))\right)\right].$
-This formulation preserves the supervised learning structure of our original objective while incorporating relative comparisons between sampled outputs, yielding richer and more informative learning signals [^1] A rule-based reward function is employed, assigning a reward of $1$ to correct answers and 0 to incorrect ones..
-The overall PACS framework is illustrated in Figure 2.
-3.4.1 Loss Function Analysis.
-To gain deeper insight into the learning dynamics, we analyze the loss function in Equation 9.
-When a sampled output $o_{i}$ is correct (i.e., $R(q,o_{i})=1$), the loss reduces to the term $\log(\sigma(\psi(q,o_{i};\pi_{\theta})))$.
-Minimizing the loss in this case corresponds to maximizing $\psi(q,o_{i};\pi_{\theta})$, which in turn requires increasing the estimated reward proxy $\hat{r}(q,o_{i};\pi_{\theta})$.
-Since $\hat{r}$ is defined as a log-probability ratio, this implies increasing $\pi_{\theta}(o_{i}|q)$, i.e., assigning higher probability to generating the correct response.
-In this way, the model learns to favor correct outputs.
-Conversely, when a sampled output $o_{i}$ is incorrect (i.e., $R(q,o_{i})=0$), the loss reduces to the term $\log(1-\sigma(\psi(q,o_{i};\pi_{\theta})))$.
-Minimizing the loss requires decreasing $\psi(q,o_{i};\pi_{\theta})$, which again corresponds to lowering the reward proxy $\hat{r}(q,o_{i};\pi_{\theta})$.
-This is achieved by reducing $\pi_{\theta}(o_{i}|q)$, discouraging the generation of incorrect responses.
+Loss Function Analysis.
+For a correct output $o_{i}$ (i.e., $R(q,o_{i})=1$), the loss reduces to $\log(\sigma(\psi(q,o_{i};\pi_{\theta})))$.
+Minimizing this term maximizes $\psi(q,o_{i};\pi_{\theta})$, thereby increases the reward proxy $\hat{r}(q,o_{i};\pi_{\theta})$ and, equivalently, the probability $\pi_{\theta}(o_{i}|q)$ of generating the correct output.
+Conversely, for an incorrect output $o_{i}$ (i.e., $R(q,o_{i})=0$), the loss becomes $\log(1-\sigma(\psi(q,o_{i};\pi_{\theta})))$.
+Minimizing it decreases $\psi(q,o_{i};\pi_{\theta})$, which in turn lowers $\hat{r}(q,o_{i};\pi_{\theta})$ and reduces the probability $\pi_{\theta}(o_{i}|q)$ of generating the incorrect output.
 Thus, the model learns not only to generate correct outputs, but also to avoid producing incorrect ones.
-3.4.2 Handling Data Imbalance.
-To address the potential distributional imbalance in generated outputs $o_{i}$ for $q$, we adopt the class imbalance treatment methodology proposed by (King & Zeng, 2001), whereby differential weights are assigned to correct and incorrect samples respectively.
-Specifically, when the proportion of correct and incorrect samples in the training data exhibits imbalance, we mitigate the adverse effects of sample distribution bias on model performance by adjusting weight parameters to balance the model’s attention across different categorical samples.
-MATH 500 ($\text{pass@}k$) AMC 23 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 56.66 67.64 75.75 81.91 86.87 90.60 30.20 66.16 83.88 90.42 94.85 97.10
-PPO 65.48 71.94 77.51 82.14 86.02 89.20 38.57 66.07 80.19 86.81 92.09 95.53
-GRPO 66.92 73.34 78.60 83.01 86.60 89.00 37.01 68.71 82.61 87.72 92.39 96.44
-PACS 67.31 74.75 80.60 85.03 88.22 90.40 39.63 68.23 82.45 88.23 92.70 96.18
-AIME 2024 ($\text{pass@}k$) AIME 2025 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 4.92 15.66 27.21 33.45 40.95 49.68 1.73 10.62 23.45 30.69 39.09 47.38
-PPO 7.55 19.70 30.27 35.69 42.08 50.81 1.40 8.99 21.11 28.07 35.52 42.51
-GRPO 8.14 19.13 26.06 31.65 37.75 42.76 2.62 7.07 14.07 20.18 28.37 36.58
-PACS 7.98 21.49 30.10 35.87 42.25 48.32 2.94 11.32 22.91 30.50 39.07 47.18
-Table 1: Results of Qwen2.5-3B trained with PPO, GRPO and PACS.
+Model MATH 500 AMC 23 AIME 2024 AIME 2025 BeyondAIME Avg.
+Closed-source Models
+Claude Opus 4.1 - 87.5 - 78.0 - -
+Gemini 2.5 pro 96.7 100.0 88.7 86.7 58.8 86.18
+Open-Source Model
+Qwen3-4B 91.01 82.30 46.48 34.71 17.33 54.37
+Qwen3-8B 89.94 81.74 46.98 33.83 16.28 53.75
+Qwen3-14B 92.07 84.88 49.06 36.22 18.59 56.16
+Deepthought-8B 45.07 16.58 1.82 0.44 0.76 12.93
+Eurus-2-7B-PRIME 82.07 63.65 18.39 13.88 6.67 36.93
+LIMO 91.00 80.61 40.86 31.82 15.98 52.05
+Marco-o1 70.48 45.12 9.22 6.95 2.77 26.91
+OpenThinker3-7B 88.90 72.44 36.95 26.98 12.33 47.52
+Sky-T1-7B 85.38 69.59 20.86 20.47 8.54 40.97
+PACS-4B 94.80 90.45 55.10 45.63 27.16 62.63
+$\Delta(\text{Qwen3-4B})$ $\uparrow$3.79 $\uparrow$8.15 $\uparrow$8.62 $\uparrow$10.92 $\uparrow$9.83 $\uparrow$8.26
+PACS-8B 95.09 88.69 57.58 46.38 28.86 63.32
+$\Delta(\text{Qwen3-8B})$ $\uparrow$5.15 $\uparrow$6.95 $\uparrow$10.6 $\uparrow$12.55 $\uparrow$12.58 $\uparrow$9.57
+Table 1: Comparison of $\text{pass@}1$ accuracy (%) on mathematical reasoning benchmarks.
+“Avg.” indicates the average performance across all tasks.
 Bold numbers indicate the best performance.
 Underlined numbers indicate the second best.
-Due to space constraints, selected results are presented here.
-The complete results are provided in AMC23 (Table 5), AIME 2024 (Table 7) and AIME 2025 (Table 9) in the Appendix.
+The values in green represent the performance gain ($\Delta$) over the model.
+Overall, this formulation preserves the supervised learning structure of our original objective while incorporating relative comparisons between sampled outputs, yielding more informative learning signals [^1] A rule-based reward function is employed, assigning a reward of $1$ to correct answers and 0 to incorrect ones..
+The overall PACS framework is illustrated in fig.
+2.
+Mitigating Entropy Collapse.
+With the score function instantiated in a policy-dependent form, the gradient of the loss can be expressed as an advantage-weighted policy gradient (Appendix A.1).
+Unlike conventional RL, where positive rewards can push the policy toward overly deterministic behavior, our gradient (eq.
+5) is modulated by the residual $R(q,o)-\sigma(\psi(q,o;\pi_{\theta}))$.
+As predictions approach the ground truth, the residual vanishes, naturally attenuating updates and preventing over-concentration of probability mass, thereby preserving a diverse set of responses.
+We provide a formal proof in Appendix A.2 showing that this mechanism implicitly imposes a structural constraint that regularizes the policy towards a Gibbs distribution, fundamentally preventing mode collapse.
+3.5 Practical Considerations: Handling Data Imbalance
+To address the potential distributional imbalance in generated outputs $o_{i}$ for $q$, we adopt the class imbalance treatment methodology proposed by (King and Zeng, 2001), whereby differential weights are assigned to correct and incorrect samples respectively.
+Specifically, when the proportion of correct and incorrect samples in the training data exhibits imbalance, we mitigate the adverse effects of sample distribution bias on model performance by adjusting weight parameters to balance the model’s attention across different categorical samples.
 4 Experiments
@@
 Datasets & Models.
-This study uniformly adopts the DeepScaleR (Luo et al., 2025) dataset as the training corpus, which constitutes a high-quality mathematical problem-solving collection comprising approximately 40,000 unique mathematical problem-answer pairs.
-To comprehensively evaluate the performance of methods, we conduct extensive evaluations on four representative benchmarks: MATH 500, AMC23, AIME 2024, and AIME 2025.
-The experiments utilize Qwen2.5-3B and Qwen2.5-7B (Qwen et al., 2025) as base models, enabling us to assess PACS’s performance across various model scales.
+We utilize the DeepScaleR dataset as the training corpus (Luo et al., 2025), which constitutes a high-quality mathematical problem-solving collection.
+For evaluation, we conduct extensive evaluations on five representative benchmarks: MATH 500, AMC 23, AIME 2024, AIME 2025 and BeyondAIME.
+The experiments utilize Qwen3-4B and Qwen3-8B (Qwen et al., 2025) as base models, assessing PACS across different model scales.
 Baselines & Hyperparameters.
-To evaluate the effectiveness of our approach, we compare the performance of PACS with representative RL algorithms: PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024).
-The training objectives of these algorithms can be found in Appendix.
+To comprehensively evaluate the effectiveness of PACS, we compare PACS with a diverse set of baselines categorized into three groups: 1).
+Closed-source models: including state-of-the art models such as Claude Opus 4.1 and Gemini 2.5 Pro (Comanici et al., 2025), whose results are directly cited from prior works (Team et al., 2025a; Li et al., 2025); 2).
+Open-source models: including recent models utilizing advanced post-training of RL techniques, such as Qwen3 models (Yang et al., 2025), Deepthought-8B (Ruliad, 2024), Eurus-2-7B-PRIME (Cui et al., 2025), LIMO (Ye et al., 2025), Marco-o1 (Zhao et al., 2024), OpenThinker3-7B (Guha et al., 2025) and Sky-T1-7B (Team, 2025a); and 3).
+Representative RL algorithms: we implement and train PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) on the same base models to ensure a fair comparison.
+The training objectives of these algorithms can be found in Appendix B.
 We also report the performance of base models for reference.
 The train batch size is set to 1,024.
-The learning rate of actor model is set to 1e-6.
-For each query, $8$ responses are sampled.
+For each query during training, $8$ responses are sampled.
 For PACS, $\beta$ is set to 1.0.
-During inference, we configure the sampling parameters with a temperature of 0.6 and top-$p$ of 0.96.
-For Qwen2.5-3B, we conduct training for 140 steps, while for Qwen2.5-7B, we extend the training to 300 steps.
-Prompt template and detailed hyperparameter configuration can be found in the Appendix.
+During inference, we configure the sampling parameters with a temperature of 0.6.
+Prompt template and detailed hyperparameter configuration can be found in the Appendix C.
 Hardware.
 All experiments are conducted on NVIDIA H200 GPUs.
-Training employs verl (Sheng et al., 2024), an efficient reinforcement learning library specifically designed for large language models.
+We employ verl (Sheng et al., 2024), an reinforcement learning library specifically designed for LLMs.
 For inference, we utilize vllm (Kwon et al., 2023).
-Evaluation Metrics.
-To evaluate the model’s capability to generate diverse correct solutions while mitigating the randomness bias that may arise from single sampling, we employ the $\text{pass@}k$ metric for assessment.
-$\text{pass@}k$ represents the probability that at least one correct answer exists among $k$ generated candidate solutions.
-$\text{pass@}k$ also enables performance evaluation under different computational budgets by adjusting the value of $k$.
-However, directly calculating $\text{pass@}k$ with only $k$ samples per instance often results in high variance.
-To obtain more stable evaluation results, we generate $n$ samples for each problem (where $n\geq k$), count the number of correct samples $c$, and calculate the unbiased estimator (Chen et al., 2021):
-$\text{pass@}k=\mathbb{E}_{x\sim D}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$]. (10)
-Our sampling strategy adapts to the scale of dataset.
-For the MATH 500 dataset, given its scale of 500 problems, the sampling size is set to $n=32$ for each problem and model performance is evaluated across $k\in\{1,2,4,8,16,32\}$.
-Conversely, for smaller test sets (AMC23, AIME 2024, and AIME 2025), which contain fewer problems but with high-quality content, larger sampling scale is employed to obtain more stable and reliable statistical results.
-Specifically, we generate 512 candidate solutions for each problem ($n=512$) and conduct evaluation across a broader range of $k$ where $k\in\{1,2,4,8,16,32,64,128,256\}$.
-MATH 500 ($\text{pass@}k$) AMC 23 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 61.55 73.90 82.09 87.48 90.93 92.80 35.72 70.21 85.53 91.19 95.89 99.18
-PPO 76.56 80.74 84.15 86.70 88.60 90.20 51.22 72.80 81.92 85.71 88.67 90.92
-GRPO 78.29 82.08 84.70 86.60 88.26 90.00 57.22 72.20 80.72 86.14 90.69 93.34
-PACS 78.97 84.39 88.18 90.68 92.33 93.60 57.83 78.17 88.22 91.95 95.51 98.51
-    - w/o weight 77.84 83.74 87.68 90.30 92.06 93.20 54.51 75.15 83.77 86.89 88.97 90.62
-AIME 2024 ($\text{pass@}k$) AIME 2025 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 7.26 23.46 36.41 43.77 52.66 62.55 3.35 16.04 28.56 34.99 42.07 50.11
-PPO 15.56 28.50 38.71 41.50 43.85 46.46 9.61 24.89 32.32 34.94 38.27 42.91
-GRPO 14.10 25.37 34.78 38.50 41.66 45.42 8.36 21.63 33.26 38.12 42.33 46.45
-PACS 15.80 29.75 43.90 51.21 56.54 59.78 13.33 29.65 42.88 49.45 54.48 58.22
-    - w/o weight 15.72 24.57 32.58 38.77 46.16 53.42 10.44 22.86 31.03 36.59 42.39 47.67
-Table 2: Results of Qwen2.5-7B trained with PPO, GRPO and PACS.
+[…]
+AMC 23 ($\text{pass@}k$) AIME 2024 ($\text{pass@}k$)
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+Base 81.74 90.53 92.31 94.00 95.65 96.86 46.98 59.78 65.10 69.28 72.12 74.99
+PPO 86.27 92.36 93.91 95.46 96.70 98.13 54.61 71.46 75.86 79.11 81.75 83.12
+GRPO 87.83 93.17 94.71 95.92 97.91 99.69 54.56 74.52 80.08 82.42 83.21 83.33
+PACS 88.69 94.28 95.36 96.03 96.72 97.35 57.58 76.24 81.21 83.30 84.14 85.00
+   - w/o weight 89.63 94.80 95.53 96.02 96.50 97.32 58.93 76.64 81.13 83.21 83.59 84.73
+AIME 2025 ($\text{pass@}k$) BeyondAIME ($\text{pass@}k$)
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+Base 33.83 45.86 51.19 56.12 60.54 65.74 16.28 23.00 26.75 31.36 36.65 41.74
+PPO 42.42 55.67 61.70 68.25 73.99 77.43 23.38 33.60 38.83 44.10 49.11 53.80
+GRPO 45.29 58.80 64.04 69.12 73.84 77.39 25.79 37.61 43.08 48.40 53.32 57.64
+PACS 46.38 61.93 67.78 72.20 75.71 79.12 28.86 43.15 49.58 55.39 60.19 64.02
+   - w/o weight 46.15 60.07 66.29 71.98 76.10 78.23 28.52 40.43 45.55 50.10 54.35 58.97
+Table 2: Results of Qwen3-8B trained with PPO, GRPO and PACS.
 Bold numbers indicate the best performance.
 Underlined numbers indicate the second best.
-Due to space constraints, selected results are presented here.
-The complete results are provided in AMC23 (Table 6), AIME 2024 (Table 8) and AIME 2025 (Table 10) in the Appendix.
+Due to space constraints, results on MATH 500 are shown in Table 8.
+4.2 Main Results
+Comparative Analysis with Existing Models.
+table 1 shows that PACS significantly enhances the reasoning capabilities of base models across all benchmarks.
+Specifically, PACS-8B achieves an average $\text{pass@}1$ of 63.32%, surpassing Qwen3-8B by 9.57 points and outperforming strong baselines like LIMO (52.05%).
+Notably, PACS exhibits exceptional parameter efficiency: PACS-4B (62.63%) not only beats 7B-scale competitors but also outperforms the significantly larger Qwen3-14B on challenging tasks like AIME 2025 (45.63% vs.
+36.22%).
+The gains are particularly pronounced on complex benchmarks; for instance, on AIME 2025 and BeyondAIME, PACS-8B improves upon the base model by over 12 points, effectively unlocking deep chain-of-thought capabilities.
+Superior Performance against RL Baselines.
+Comparing algorithmic efficacy in tables 2 and 6, PACS consistently outperforms PPO and establishes a decisive advantage over GRPO on complex reasoning tasks.
+This superiority is most evident on the hardest benchmark, BeyondAIME.
+Here, PACS-8B achieves a $\text{pass@}64$ of 64.02%, surpassing GRPO (57.64%) by over 6 points, while PACS-4B similarly improves the metric from 57.58% to 62.10%.
+Furthermore, PACS-8B maintains a comprehensive lead over GRPO on AIME 2024 and 2025 in both single- and multi-sample settings.
+These results indicate that PACS’s supervised learning formulation, which implicitly couples actor and critic updates, provides more effective optimization signals than pure RL approaches for navigating complex solution spaces.
+Training Dynamics and Performance Analysis.
+To further investigate the optimization efficiency, we visualize the evaluation trajectories of Qwen3-8B throughout the training process in Figure 3.
+The curves report $\text{pass@}1$ on benchmarks, evaluated every 20 training steps.
+The results highlight PACS’s robustness across diverse problem distributions.
+On benchmarks like AMC 23 and the highly challenging BeyondAIME, PACS exhibits high training efficiency, rapidly establishing a decisive performance lead over RL baselines.
+Notably, on AIME 2025, the performance gap narrows as GRPO also demonstrates competitive learning trajectories; yet, PACS remains resilient, matching the strong baseline’s pace and securing the best final performance.
+Figure 3: Training dynamics of Qwen3-8B. The curves illustrate the evolution of $\text{pass}@1$ on math benchmarks throughout the training process.
+5 Ablation Study
+5.1 Different $\beta$
 Figure 4: Performance analysis of PACS with varying $\beta$.
-The 3D heatmaps show pass@k scores for different combinations of $\beta$ values (0.1, 0.5, 1, 2, 10) and $k$ values on MATH-500, AMC23, AIME-2024, and AIME-2025 datasets.
-4.2 Main Results
-Table 1 and Table 2 compare the overall results of PACS with different baselines.
-Experimental results on the Qwen2.5-3B model demonstrate that PACS exhibits significant performance advantages across multiple mathematical reasoning tasks.
-On the MATH 500 dataset, PACS achieves optimal performance across all $k$-value configurations, particularly attaining a pass rate of 67.31% at $\text{pass@}1$, representing an improvement of 10.65 points over the base model and surpassing both PPO and GRPO.
-In the AMC 23 task, PACS similarly demonstrates exceptional performance.
-Notably, on the more challenging AIME 2024 and AIME 2025, PACS maintains its leading position across most $k$-value settings, demonstrating the algorithm’s stability and effectiveness in complex mathematical reasoning tasks.
-On the larger-scale Qwen2.5-7B model, the advantages of the PACS algorithm become even more pronounced, further validating its generalization capability across different model scales.
-On the MATH 500 dataset, PACS achieves superior performance across all evaluation metrics.
-The performance of PACS is especially remarkable on the AIME series tasks: in AIME 2024, it achieves a pass rate of 59.78% at $\text{pass@}256$, representing improvements of 13.32 and 14.36 points compared to PPO and GRPO, respectively; in AIME 2025, the $\text{pass@}256$ metric shows improvements of 15.31 and 11.77 points compared to PPO and GRPO, respectively.
-These results conclusively demonstrate PACS’s exceptional performance in handling high-difficulty mathematical reasoning problems and its significant advantages over reinforcement learning methods PPO and GRPO.
-4.3 Ablation Study
-Different $\beta$.
-To validate the impact of $\beta$ on the model’s performance in Equation 8, we design and conduct ablation experiments.
-Models are trained with different $\beta\in\{0.1,0.5,1,2,10\}$ and their performance is evaluated in Figure 3.
-The results demonstrate that model achieves optimal performance when $\beta=1$, yielding higher $\text{pass@}k$.
-Further analysis reveals that the AIME 2024 and AIME 2025 datasets exhibit heightened sensitivity to variations in $\beta$.
-This suggests that for tasks of higher complexity and difficulty, $\beta$ constitutes a critical step in achieving optimal performance.
-In contrast, the AMC23 and MATH 500 datasets demonstrate greater robustness across different $\beta$, with minor fluctuations in model performance.
-This phenomenon can be attributed to the relatively simpler nature of problems in these datasets, enabling the model to maintain stable performance across a broader range of hyperparameters.
-Different Advantage Estimators.
-To investigate the influence of advantage estimators on PACS’s efficacy, we extend our evaluation beyond the default RLOO to include GRPO (Shao et al., 2024) and Dr.
-GRPO (Liu et al., 2025b) as shown in Appendix C.4.
-Our findings indicate that the PACS with RLOO exhibits superior overall performance, a lead that becomes more significant at higher $\text{pass@}k$ metrics.
-Notably, whereas RLOO and Dr.
-GRPO yield comparable results on datasets such as MATH 500 and AMC23, RLOO’s superiority is markedly evident on the more demanding AIME benchmark.
-We posit that this performance gap stems from a subtle algorithmic divergence between RLOO and Dr.
-GRPO, which is magnified under more challenging conditions.
-By contrasting a single sample with the cohort average, RLOO’s leave-one-out mechanism yields more robust and precise signals for credit assignment.
-Effect of Weight Integration.
-To validate the effectiveness of the weight mechanism in PACS, we conduct ablation experiments by comparing the performance of PACS with and without the weight component.
-The results presented in Table 2 clearly demonstrate the positive impact of the weight mechanism.
-PACS consistently outperforms its weight-free variant (w/o weight).
-The benefits are particularly pronounced on challenging AIME datasets, where the weight component proves crucial for complex mathematical reasoning tasks, achieving substantial performance improvements especially at higher sampling rates (e.g., 6.36% point improvement on AIME 2024 at $\text{pass@}32$ of Qwen2.5-7B).
-4.4 Training Dynamics
-Figure 4: Training dynamics of Qwen2.5-7B on the DeepScaleR dataset across different optimization methods (PPO, GRPO, PACS, and PACS w/o weight).
-The plots illustrate the evolution of three key metrics during training: (a) entropy loss, (b) gradient norm, and (c) mean response length over training steps.
-Figure 4 compares training dynamics of four methods applied to Qwen2.5-7B.
-PPO and GRPO maintain consistently low entropy loss, resulting in entropy collapse and conservative outputs.
-But PACS shows a non-monotonic pattern: entropy increases to about 1,700 at step 220, then stabilizes at 1,200-1,300, reflecting a superior exploration-exploitation balance.
-PACS w/o weight demonstrates intermediate behavior.
-Gradient norms reveal similar patterns.
-While PPO and GRPO maintain stable gradient norms throughout training, PACS peaks at 40 with sustained elevation, enabling aggressive parameter updates and broader exploration.
-This sustained gradient elevation in PACS, contrasting with the stable but lower gradient norms in baseline methods, demonstrates superior training efficiency.
-The consistently higher gradient activity indicates that PACS continues making meaningful parameter updates throughout the training process, while the stable but modest gradients in PPO and GRPO suggest more conservative optimization that may lead to premature convergence.
-Response lengths show PACS generating the most detailed responses.
-PACS achieves superior exploration capability and response quality through moderate entropy maintenance and dynamic gradient updates, providing novel optimization insights compared to other methods.
+The 3D heatmaps show $\text{pass@}k$ scores for different combinations of $\beta$ values (0.1, 0.5, 1, 2, 10) and $k$ values on AMC23, AIME-2024, AIME-2025 and BeyondAIME.
+To validate the impact of $\beta$ on the model’s performance in eq.
+7, we design and conduct ablation experiments.
+Models are trained with different $\beta\in\{0.1,0.5,1,2,10\}$ and their performance is evaluated in fig.
+4.
+The results demonstrate that PACS is highly robust to variations in $\beta$.
+While extreme values (e.g., $\beta=10$) maintain competitive performance on easier tasks like AMC 23, a distinct ”sweet pot” emerges in $\beta\in[0.5,1.0]$ for maximizing $\text{pass@}1$ on complex reasoning tasks.
+Given that $\text{pass@}1$ is often the primary objective for reasoning models, we select $\beta=1$ as the optimal setting, effectively balancing the reward signal strength to achieve superior reasoning performance.
+5.2 Different Score Function
+To validate the effectiveness of our scoring mechanism, we compare the RLOO score function against GRPO (Shao et al., 2024) and Dr.
+GRPO (Liu et al., 2025c).
+Detailed results are shown in Appendix D.3.
+Our analysis reveals that PACS with RLOO generally yields superior performance, particularly on complex reasoning tasks.
+For instance, on BeyondAIME benchmark, PACS achieves a $\text{pass@}1$ of 27.16%, consistently outperforming both GRPO (26.41%) and Dr.
+GRPO (26.67%).
+We attribute this robustness to the leave-one-out mechanism of RLOO, which provides unbiased advantage estimates with lower variance compared to the group-mean baselines used in GRPO variants.
+5.3 Effect of Weight Integration
+To validate the effectiveness of the weight mechanism in PACS, we conduct ablation experiments by comparing the performance of PACS with and without the weight component as shown in tables 2, 6, 7 and 8.
+For Qwen3-4B, PACS systematically outperforms the unweighted variant across benchmarks; notably, it achieves a pass@1 of 90.45% on AMC 23 and 55.10% on AIME 2024, surpassing the w/o weight PACS scores of 89.18% and 54.56% respectively.
+This advantage is further amplified on complex tasks with the larger Qwen3-8B model.
+While performance on simpler tasks remains comparable, PACS establishes a clear lead on the hardest benchmark, BeyondAIME.
+It not only improves the $\text{pass@}1$ accuracy to 28.86% compared to 28.52% , but more importantly, achieves a remarkable 64.02% at pass@64, exceeding the unweighted variant’s 58.97%.
+Figure 5: Exploration and Diversity Analysis.
+(a) Entropy loss dynamics for Qwen3-4B (top) and 8B (bottom).
+Unlike baselines that suffer from entropy collapse, PACS maintains higher entropy, enabling sustained exploration.(b) Centered PCA projection of correct solutions for sampled problems (ID shown in bottom-right).
+The broader semantic coverage of PACS (Blue) compared to GRPO (Red) visually confirms superior diversity.
+6 Generalization and Diversity Analysis
+In this section, we analyze the broader capabilities of PACS by first evaluating its generalization to out-of-domain tasks to rule out overfitting, and then exploring the solution diversity that underpin this mechanism.
+6.1 Generalization of Out-of-Domain Tasks
+Figure 6: Performance on out-of-domain tasks. The chart reports the evaluation scores of Qwen3-8B on LiveCodeBench and GPQA. PACS achieves substantial gains over PPO and GRPO.
+A key challenge in RLVR is preventing the policy from over-specializing to the training domain, which can lead to a degradation of general capabilities.
+To evaluate whether PACS learns fundamental reasoning skills that transfer across domains, we assess the Qwen3-8B model trained on the DeepScaleR dataset (math) on two out-of-domain benchmarks: LiveCodeBench (Jain et al., 2024) [^2] Evaluations are performed on the LiveCodeBench dynamic subset spanning the period from 7/1/2024, through 1/1/2025.
+and GPQA (Rein et al., 2023).
+For these assessments, we utilize the toolkit provided by GLM-4.5, adhering to the default settings for all experimental configurations.
+As shown in Figure 6, PACS demonstrates superior cross-domain generalization compared to PPO and GRPO.
+On LiveCodeBench, PACS achieves 45.96, significantly outperforming the Vanilla model (31.06) as well as PPO (35.87) and GRPO (38.01), which suggests that the reasoning chains optimized by PACS effectively transfer to code logic generation while baseline RL methods yield diminishing returns.
+This advantage is further corroborated on GPQA, where PACS attains 52.46, surpassing both GRPO (48.55) and PPO (48.48).
+These results indicate that PACS does not merely memorize domain-specific templates; instead, by treating outcome rewards as supervised signals within a stable optimization framework, it enhances the model’s intrinsic reasoning capabilities.
+6.2 Exploration and Diversity
+Model Method AIME2024 AIME2025
+Qwen3-4B PACS 0.0278 0.0300
+GRPO 0.0260 0.0266
+Qwen3-8B PACS 0.0272 0.0392
+GRPO 0.0217 0.0249
+Table 3: Comparison of diversity scores on AIME 2024 and AIME 2025. PACS consistently exhibits higher diversity than GRPO across different model scales.
+We observe that PACS effectively mitigates the entropy collapse issue commonly found in RL baselines, leading to semantically richer and more diverse solutions as shown in fig.
+5(a).
+To quantify the diversity of the generated reasoning paths, we utilize the average pairwise consine distance between the embeddings of responses.
+A higher diversity score ($S_{div}$) indicates greater semantic variation.
+The definition of $S_{div}$ is detailed in Appendix E.
+We assess the response diversity on two representative benchmarks: AIME 2024 and AIME 2025.
+We specifically compare PACS with GRPO using the Qwen3-4B and Qwen3-8B.
+Note that we compute $S_{div}$ on correct responses, as our goal is to achieve diverse yet valid solutions.
+As shown in table 3, PACS consistently surpasses GRPO in diversity (0.0392 vs.
+0.0249).
+While the rigid nature of mathematical solutions limits absolute scores, PACS achieves a substantial relative improvement, effectively mitigating mode collapse and fostering semantically distinct reasoning paths.
+To intuitively compare the diversity of reasoning paths, centered PCA approach is employed.
+We randomly sample some problems.
+fig.
+5(b) displays the 2D PCA projection of the embeddings for correct solutions generated by PACS (Blue) and GRPO (Red).
+The data is mean-centered to align the centroids of both distributions at the origin $(0,0)$.
+The shaded regions (KDE contours) and scatter points illustrate that PACS covers a significantly larger area in the semantic space, indicating a more diverse set of reasoning paths compared to the more concentrated distribution of GRPO.
 7 Conclusion
-In this work, we propose PACS, a novel RLVR algorithm that achieves implicit actor-critic coupling via supervised learning.
-By treating outcome rewards as supervised targets and parameterizing the score function with policy log probabilities, PACS addresses the sparse reward and training instability challenges in existing RL-based RLVR methods.
-Extensive experiments on four mathematical benchmark datasets demonstrate that PACS significantly outperforms strong baselines including PPO and GRPO, achieving superior task performance while maintaining healthier policy entropy for sustained exploration.
-These findings suggest that supervised learning-based implicit actor-critic coupling represents a promising paradigm for advancing RLVR.
-Appendix A Baselines
-A.1 Proximal Policy Optimization Algorithms (PPO)
+In this work, we propose PACS, a novel RLVR algorithm that achieves implicit actor-critic coupling via a supervised learning framework.
+By treating outcome rewards as supervised targets and parameterizing the score function with policy log probabilities, PACS mitigates the sparse reward and training instability challenges which are inherent to existing RL-based methods.
+Extensive experiments on various benchmarks demonstrate that PACS significantly outperforms strong baselines including PPO and GRPO, achieving superior task performance while maintaining healthier policy entropy for sustained exploration.
+These results highlight PACS as a promising approach for advancing RLVR.
+Appendix A Proofs and Theoretical Discussions
+In this section, we provide a rigorous theoretical analysis of the PACS framework.
+We first derive the exact gradient update of the proposed objective function, demonstrating that it inherently recovers an advantage-weighted policy gradient.
+Subsequently, we analyze the global optimality conditions to explain why PACS maintains solution diversity and resists entropy collapse, a property achieved through implicit regularization rather than explicit penalty terms.
+A.1 Derivation of the PACS Gradient Update
+We first formally establish the gradient dynamics of PACS.
+We demonstrate that minimizing the supervised cross-entropy loss over the score function $\psi(q,o;\pi_{\theta})$ results in a policy update mathematically equivalent to the policy gradient.
+Specifically, the effective advantage function in this update is dynamically composed of the cross-entropy loss term and the reward prediction error.
+Proposition A.1 (Gradient Derivation).
+The gradient of the PACS objective function $\mathcal{L}_{\text{PACS}}(\theta)$ with respect to the policy parameters $\theta$ can be formulated as a policy gradient update weighted by an effective advantage function derived from the prediction error and the cross-entropy loss.
+Proof.
+Consider the PACS objective defined as the expected binary cross-entropy loss:
+$\mathcal{L}_{\text{PACS}}(\theta)=-\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(\cdot|q)}\left[\ell_{\text{BCE}}(R(q,o),\sigma(\psi_{\theta}(q,o)))\right]$
+[…]
+To compute the gradient $\nabla_{\theta}\mathcal{L}_{\text{PACS}}(\theta)$, we apply the log-derivative trick to the expectation term and the chain rule to the integrand:
+$\nabla_{\theta}\mathcal{L}_{\text{PACS}}=-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\cdot\ell_{\text{BCE}}(q,o;\pi_{\theta})+\nabla_{\theta}\ell_{\text{BCE}}(q,o;\pi_{\theta})\right]$
+For the second term, $\nabla_{\theta}\ell_{\text{BCE}}(q,o;\pi_{\theta})$, we apply the chain rule with respect to the score $\psi_{\theta}$.
+Noting that the derivative of the sigmoid function is $\sigma^{\prime}=\sigma(1-\sigma)$ and $\frac{\partial\ell_{\text{BCE}}}{\partial\hat{y}}=\frac{\hat{y}-y}{\hat{y}(1-\hat{y})}$, we obtain:
+$\nabla_{\theta}\ell_{\text{BCE}}$ $=\frac{\partial\ell_{\text{BCE}}}{\partial\sigma}\frac{\partial\sigma}{\partial\psi}\nabla_{\theta}\psi_{\theta}$
+$=\frac{R-\sigma(\psi_{\theta})}{\sigma(\psi_{\theta})(1-\sigma(\psi_{\theta}))}\cdot\sigma(\psi_{\theta})(1-\sigma(\psi_{\theta}))\cdot\nabla_{\theta}\psi_{\theta}$
+$=(R(q,o)-\sigma(\psi_{\theta}))\nabla_{\theta}\psi_{\theta}$
+Given the definition of $\psi_{\theta}$, its gradient is $\nabla_{\theta}\psi_{\theta}=\beta\nabla_{\theta}\log\pi_{\theta}(o|q)$ (since $\pi_{\text{ref}}$ is fixed). Substituting this back yields:
+$\nabla_{\theta}\ell_{\text{BCE}}=\beta(R(q,o)-\sigma(\psi_{\theta}))\nabla_{\theta}\log\pi_{\theta}(o|q)$
+Combining both terms, we obtain the final gradient expression:
+$\nabla_{\theta}\mathcal{L}_{\text{PACS}}=-\mathbb{E}_{q,o\sim\pi_{\theta}(\cdot|q)}\left[\underbrace{\left(\ell_{\text{BCE}}(q,o;\pi_{\theta})+\beta(R(q,o)-\sigma(\psi_{\theta}))\right)}_{\text{Effective Advantage }A_{\text{PACS}}}\nabla_{\theta}\log\pi_{\theta}(o|q)\right]$
+This result confirms that PACS performs a valid policy gradient update.
+The effective advantage $A_{\text{PACS}}$ naturally incorporates a “critic” signal (the residual $R-\sigma(\psi_{\theta})$) which vanishes as the model’s prediction aligns with the ground truth, providing a variance-reduced update signal compared to standard policy gradients with statistical baselines.
+∎
+A.2 Implicit Regularization and Robustness to Entropy Collapse
+A key empirical finding of this work is that PACS effectively prevents entropy collapse (the degeneration of the policy into a deterministic distribution), a common failure mode in standard RLVR methods like GRPO when explicit KL penalties are omitted.
+In this section, we provide a theoretical justification for this robustness.
+We prove that the structural parameterization of PACS imposes an implicit regularization that constrains the optimal policy to a Gibbs distribution, thereby preserving exploration.
+Proposition A.2 (Implicit Gibbs Regularization).
+Assuming sufficient model capacity, the global minimum of the PACS objective function is achieved if and only if the policy $\pi_{\theta}$ follows a Gibbs distribution anchored to the reference policy $\pi_{\text{ref}}$.
+This stands in contrast to standard unregularized RL objectives, where the global minimum is a deterministic Dirac distribution.
+Proof.
+We first analyze the optimality condition for the PACS objective.
+Since $\mathcal{L}_{\text{PACS}}$ is a strictly convex binary cross-entropy loss with respect to the logits $\psi_{\theta}$, for any query-response pair $(q,o)$, the loss is minimized when the predicted probability matches the expected target reward:
+$\sigma(\psi_{\theta}^{*}(q,o))=\mathbb{E}_{\text{data}}[R(q,o)]$
+In the context of RLVR with deterministic ground-truth verification, this implies the model attempts to predict the correctness of $o$.
+Inverting the sigmoid function, the optimal score function $\psi_{\theta}^{*}$ must satisfy:
+$\psi_{\theta}^{*}(q,o)=\sigma^{-1}(\mathbb{E}[R(q,o)])$
+Crucially, in PACS, the score function is not a free parameter but is structurally coupled to the policy via the definition $\psi_{\theta}(q,o)=\beta(\log\pi_{\theta}(o|q)-\log\pi_{\text{ref}}(o|q))$.
+Substituting this definition into the optimality condition yields:
+$\beta\log\frac{\pi_{\theta}^{*}(o|q)}{\pi_{\text{ref}}(o|q)}=\sigma^{-1}(\mathbb{E}[R(q,o)])$
+Exponentiating both sides and rearranging terms, we derive the form of the optimal policy $\pi_{\theta}^{*}$:
+$\pi_{\theta}^{*}(o|q)=\pi_{\text{ref}}(o|q)\exp\left(\frac{\sigma^{-1}(\mathbb{E}[R(q,o)])}{\beta}\right)$
+Comparison with Standard RL: Consider a standard RL objective $J=\mathbb{E}_{o\sim\pi}[R(o)]$ (often optimized by baselines without explicit KL).
+The global maximum of $J$ occurs at a deterministic Dirac delta distribution: $\pi_{\text{std}}^{*}=\delta(o-\operatorname*{arg\,max}_{o}R(o))$, which represents total entropy collapse.
+In contrast, the optimal policy for PACS derived above is a Gibbs distribution.
+The term $\pi_{\text{ref}}(o|q)$ acts as a base measure, ensuring that the optimized policy retains the support and distributional characteristics of the pre-trained model.
+Even without an explicit KL loss term in the training objective, the specific parameterization of $\psi_{\theta}$ enforces this structural constraint.
+This guarantees that PACS inherently balances reward maximization with adherence to the reference distribution, theoretically explaining the superior diversity and resistance to mode collapse observed in our experiments.
+∎
+Appendix B Baselines
+B.1 Proximal Policy Optimization Algorithms (PPO)
 The objective function of PPO (Schulman et al., 2017) is formulated as follows:
-$\mathcal{J}_{\text{PPO}}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta_{\text{old}}}(\cdot|q)}\frac{1}{|o|}\sum_{t=1}^{|o|}\left\{\min\left[\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}A_{t},\text{clip}\left(\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})},1-\varepsilon,1+\varepsilon\right)A_{t}\right]\right\}$
+$\mathcal{J}_{\text{PPO}}(\theta)=$ $\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta_{\text{old}}}(\cdot|q)}$
+$\frac{1}{|o|}\sum_{t=1}^{|o|}\left\{\min\left[\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}A_{t},\text{clip}\left(\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})},1-\varepsilon,1+\varepsilon\right)A_{t}\right]\right\}$
 where $A$ represents the advantage function estimated using Generalized Advantage Estimation (GAE), and $\varepsilon$ denotes the clipping hyperparameter that stabilizes the training process in PPO.
-A.2 Group Relative Policy Optimization (GRPO)
+B.2 Group Relative Policy Optimization (GRPO)
 The objective function of GRPO (Shao et al., 2024) is defined as:
-$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|q)}$
-$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}A_{i,t},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)A_{i,t}\right]-\beta D_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\right\}$
+$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\\ \Bigg\{\min\left[\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}A_{i,t},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)A_{i,t}\right]-\beta D_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\Bigg\}$ (9)
 where the advantage function is computed as:
@@
 GRPO (Liu et al., 2025b), the KL divergence term may be optionally omitted from the original GRPO objective (i.e., setting $\beta=0$), as empirical evidence suggests it may not be essential for performance.
-Appendix B Hyperparameters Settings
+Appendix C Hyperparameters Settings
 To ensure Experimental efficiency and effectiveness, specialized optimized frameworks are used for training and inference phases. This section details the key hyperparameters used in each phase.
-B.1 Template
+C.1 Template
 […]
@@
 <|im_start|>assistant
-B.2 Training Hyperparameters
+C.2 Training Hyperparameters
 During the model training phase, the verl (Sheng et al., 2024) framework is employed, which is a powerful toolkit designed for large-scale model training.
@@
 Train batch size 1,024
-Max prompt length 512
-Max response length 4,096
+Max prompt length 1024
+Max response length 8,192
 Filter overlong prompts True
 Mini batch size 256
-Micro batch size per GPU 16 for Qwen2.5-3B, 8 for Qwen2.5-7B
+Micro batch size per GPU 8
 Learning rate $1\times 10^{-6}$ for actor, $1\times 10^{-5}$ for critic in PPO
@@
 Use KL loss False
-Total training steps 140 for Qwen2.5-3B, 300 for Qwen2.5-7B
+Total training steps 120 for Qwen3-4B, 320 for Qwen3-8B
 Table 4: Training hyperparameters.
-B.3 Inference Hyperparameters
+C.3 Inference Hyperparameters
 During the model inference and evaluation phase, the high-performance inference serving framework vllm (Kwon et al., 2023) is employed.
@@
 GPU memory utilization 0.9
-Max tokens 4,096
+Max tokens 8,192
 Temperature 0.6
 Top-pp 0.95
-Tensor parallel size 4
+Tensor parallel size 1
 Table 5: Inference hyperparameters.
-Appendix C Detailed Experimental Results
-This section presents the comprehensive test results of each model on AMC 23, AIME 2024, and AIME 2025, along with corresponding analyses.
-C.1 AMC23
-The experimental results clearly demonstrate the superior performance of the PACS algorithm on the AMC23.
-As shown in Table 5 and Table 6, PACS achieves optimal or near-optimal performance across the majority of $\text{pass@}k$ metrics.
-The advantages of PACS are particularly pronounced at lower $k$ values; for instance, on the Qwen2.5-3B and Qwen2.5-7B models, its pass@1 scores reach 39.63 and 57.83 respectively, surpassing all baselines.
-This indicates that the proposed algorithm can more effectively enhance the model’s capability to solve problems in the first attempt or within a few trials, which holds significant value in practical applications.
-As the model scale expands from 3B to 7B, the PACS algorithm demonstrates excellent scalability, with its performance advantages becoming increasingly pronounced.
-On the Qwen2.5-7B model, PACS not only achieves leading performance in pass@1 but also comprehensively outperforms other baselines across the broad range from $k=1$ to $k=64$, showing substantial performance improvements.
-For example, at pass@8, PACS (78.17) outperforms the second-best GRPO (72.20) by nearly 6 points.
-AMC23 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 30.20 42.56 55.10 66.16 75.62 83.88 90.42 94.85 97.10
-PPO 38.57 48.65 57.97 66.07 73.25 80.19 86.81 92.09 95.53
-GRPO 37.01 46.85 58.14 68.71 76.69 82.61 87.72 92.39 96.44
-PACS 39.63 50.56 60.30 68.23 75.97 82.45 88.23 92.70 96.18
-Table 5: Results of Qwen2.5-3B on AMC23 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
-AMC23 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 35.72 49.47 61.05 70.21 78.35 85.53 91.19 95.89 99.18
-PPO 51.22 60.01 67.22 72.80 77.52 81.92 85.71 88.67 90.92
-GRPO 57.22 63.36 68.21 72.20 76.06 80.72 86.14 90.69 93.34
-PACS 57.83 65.57 71.94 78.17 83.83 88.22 91.95 95.51 98.51
-    - w/o weight 54.51 62.95 69.65 75.15 79.77 83.77 86.89 88.97 90.62
-Table 6: Results of Qwen2.5-7B on AMC23 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
-C.2 AIME 2024
-On the more challenging AIME 2024 dataset, PACS again demonstrates its effectiveness.
-On the Qwen2.5-3B model scale, different algorithms show very similar performance with intense competition, yet PACS still achieves leading results in critical intervals.
-Throughout the entire $\text{pass@}k$ curve, PACS consistently remains in the top tier.
-On the Qwen2.5-7B model, PACS achieves comprehensive leadership from $k=1$ to $k=128$, obtaining optimal results at nearly all evaluation points.
-Similar to the AMC experiments, the base model again achieves the highest score at $k=256$, which further confirms the inference that unaligned base models retain higher answer diversity and have the potential to “enumerate” correct answers through massive sampling, though this approach is highly inefficient for practical applications.
-AIME 2024 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 4.92 7.90 11.43 15.66 21.11 27.21 33.45 40.95 49.68
-PPO 7.55 10.83 14.96 19.70 24.94 30.27 35.69 42.08 50.81
-GRPO 8.14 11.45 14.79 18.13 21.74 26.06 31.65 37.75 42.76
-PACS 7.98 12.28 16.98 21.49 25.58 30.10 35.87 42.25 48.32
-Table 7: Results of Qwen2.5-3B on AIME 2024 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
-AIME 2024 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 7.26 11.66 17.17 23.46 29.94 36.41 43.77 52.66 62.55
-PPO 15.56 19.45 23.51 28.50 34.16 38.71 41.50 43.85 46.46
-GRPO 14.10 16.92 20.71 25.37 30.34 34.78 38.50 41.66 45.42
-PACS 15.80 19.93 24.39 29.75 36.29 43.90 51.21 56.54 59.78
-    - w/o weight 15.72 19.60 22.24 24.57 27.88 32.58 38.77 46.16 53.42
-Table 8: Results of Qwen2.5-7B on AIME 2024 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
-C.3 AIME 2025
-To comprehensively evaluate the performance stability and generalization capability of PACS, we conduct assessments on the latest AIME 2025 benchmark.
-This dataset represents the current highest difficulty standard for mathematical reasoning tasks, enabling more rigorous examination of algorithm performance under extremely challenging scenarios.
-On 3B-scale models, PACS demonstrates impressive comprehensive advantages, achieving optimal performance across nearly all evaluation metrics.
-PACS obtains superior results throughout the entire evaluation spectrum from pass@1 to pass@256.
-On 7B models, PACS achieves complete dominance, from pass@1 of 13.33 to pass@256 of 58.22, with every metric substantially outperforming all baseline methods.
-It is worth emphasizing that the performance gap between PACS and the second-ranked method is highly significant across multiple key metrics.
-For instance, at pass@16, PACS (36.00) surpasses GRPO (27.57) by over 8 percentage points; at pass@128, the gap reached 12 percentage points (54.48 vs 42.33).
-AIME 2025 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 1.73 3.32 6.14 10.62 16.67 23.45 30.69 39.09 47.38
-PPO 1.40 2.71 5.07 8.99 14.56 21.11 28.07 35.52 43.51
-GRPO 2.62 3.92 5.29 7.07 9.87 14.07 20.18 28.37 36.58
-PACS 2.94 4.92 7.63 11.32 16.47 22.91 30.50 39.07 47.18
-Table 9: Results of Qwen2.5-3B on AIME 2025 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
-AIME 2025 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-Base 3.35 6.12 10.40 16.04 22.32 28.56 34.99 42.07 50.11
-PPO 9.61 14.42 19.69 24.89 29.21 32.32 34.94 38.27 42.91
-GRPO 8.36 11.82 16.34 21.63 27.57 33.26 38.12 42.33 46.45
-PACS 13.33 18.19 23.73 29.65 36.00 42.88 49.45 54.48 58.22
-    - w/o weight 10.44 15.12 19.36 22.86 26.48 31.03 36.59 42.39 47.67
-Table 10: Results of Qwen2.5-7B on AIME 2025 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
-Synthesizing results across the four benchmarks—MATH 500, AMC23, AIME 2024, and AIME 2025—PACS exhibits exceptional cross-task generalization capability and performance consistency.
-Whether on the relatively straightforward MATH 500 and AMC23 or the extremely challenging AIME 2025, PACS consistently outperforms existing mainstream methods.
-This conclusively demonstrates that our proposed algorithm possesses robust adaptability and resilience, enabling outstanding performance across mathematical reasoning tasks of varying difficulty levels.
-C.4 Ablation Study
+Appendix D Detailed Experimental Results
+This section contains additional experimental results that are omitted from the main text due to space constraints.
+D.1 Results of Qwen3-4B
+Detailed performance metrics for Qwen3-4B on the AMC 23, AIME 2024, AIME 2025, and BeyondAIME benchmarks are presented below.
+AMC 23 ($\text{pass@}k$) AIME 2024 ($\text{pass@}k$)
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+Base 82.30 91.15 93.33 95.34 97.38 99.20 46.48 61.00 66.15 70.21 73.77 77.21
+PPO 86.33 93.93 95.48 96.37 97.07 97.47 52.21 70.64 77.03 81.12 82.94 83.32
+GRPO 86.95 94.69 95.98 97.08 98.44 99.70 52.84 70.71 76.31 79.97 82.05 83.13
+PACS 90.45 96.01 96.96 97.40 97.50 97.50 55.10 72.91 77.93 80.74 82.25 83.13
+   - w/o weight 89.18 94.91 95.80 96.54 97.18 97.48 54.56 72.60 78.32 81.68 83.13 83.33
+AIME 2025 ($\text{pass@}k$) BeyondAIME ($\text{pass@}k$)
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+Base 34.71 49.27 54.57 58.40 62.41 66.81 17.33 24.92 29.31 34.37 39.91 45.29
+PPO 40.34 54.87 60.52 66.36 71.58 75.24 21.82 32.97 39.06 45.46 51.28 56.01
+GRPO 40.18 54.87 61.23 67.69 73.48 77.34 24.39 37.90 44.31 49.52 53.79 57.58
+PACS 45.63 61.00 66.18 70.14 72.91 74.80 27.16 41.66 47.96 53.41 57.91 62.10
+   - w/o weight 44.90 59.10 64.08 68.11 71.37 73.09 25.89 39.11 45.28 51.18 56.53 61.12
+Table 6: Results of Qwen3-4B trained with PPO, GRPO and PACS.
+Bold numbers indicate the best performance.
+Underlined numbers indicate the second best.
+Due to space constraints, results on MATH 500 are shown in table 7.
+D.2 MATH500
+MATH500 ($\text{pass@}k$)
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+Base 91.00 95.75 96.86 97.68 98.21 98.50
+PPO 93.25 97.06 97.86 98.40 98.80 99.13
+GRPO 93.91 97.48 98.20 98.74 99.17 99.43
+PACS 94.80 97.87 98.47 98.93 99.30 99.49
+   - w/o weight 94.35 97.70 98.41 98.92 99.23 99.35
+Table 7: Results of Qwen3-4B on MATH500 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
+MATH500 ($\text{pass@}k$)
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+Base 89.94 95.20 96.54 97.54 98.27 98.75
+PPO 93.78 97.16 97.88 98.38 98.90 99.40
+GRPO 95.15 98.08 98.67 99.13 99.50 99.69
+PACS 95.09 98.09 98.71 99.16 99.50 99.72
+   - w/o weight 95.32 98.06 98.66 99.13 99.46 99.65
+Table 8: Results of Qwen3-8B on MATH500 trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
+We further analyze the performance on the MATH500, as detailed in tables 7 and 8.
+On the Qwen3-4B scale, PACS demonstrates a consistent and clear advantage over the baselines.
+Specifically, it achieves a $\text{pass@}1$ score of 94.80%, surpassing GRPO (93.91%) by roughly 0.9 points, and maintains this lead across all $k$ values, reaching 99.49% at $\text{pass@}64$.
+On the Qwen3-8B scale, performance on this benchmark approaches saturation, with base models already scoring near 90%.
+Despite this limited room for improvement, PACS remains highly competitive.
+It significantly outperforms PPO and matches the strong performance of GRPO at $\text{pass@}1$ (95.09% vs.
+95.15%), while marginally surpassing it at higher sample counts (e.g., $k\geq 4$).
+This confirms that PACS retains its stability and effectiveness even on relatively simpler tasks where the solution space is well-explored.
+D.3 Ablation of Different Advantage Estimators
 To conduct a comparative analysis of how different advantage estimators affect PACS performance, we benchmark our default RLOO method against two alternatives: GRPO (Shao et al., 2024) and Dr.
-GRPO (Liu et al., 2025b).
+GRPO (Liu et al., 2025c).
 While RLOO utilizes a leave-one-out mechanism, the advantage functions for GRPO and Dr.
@@
 MATH500 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$
-PACS 78.97 84.39 88.18 90.68 92.33 93.60
-    - GRPO 74.71 80.19 84.31 87.42 89.64 91.20
-    - Dr. GRPO 79.68 84.44 87.90 90.52 92.30 93.40
-Table 11: Results of different advantage estimators of Qwen2.5-7B on MATH 500. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+PACS 94.80 97.87 98.47 98.93 99.30 99.49
+   - GRPO 94.58 97.78 98.49 99.00 99.27 99.38
+   - Dr. GRPO 94.82 97.86 98.48 98.93 99.27 99.47
+Table 9: Results of different advantage estimators of Qwen3-4B on MATH 500. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
 AMC23 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-PACS 57.83 65.57 71.94 78.17 83.83 88.22 91.95 95.51 98.51
-    - GRPO 48.66 57.43 63.96 69.68 75.71 81.63 86.94 90.95 94.06
-    - Dr. GRPO 58.29 66.54 72.83 78.27 83.28 87.45 90.12 91.65 93.13
-Table 12: Results of different advantage estimators of Qwen2.5-7B on AMC23. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
-On the MATH 500 and AMC23 benchmarks (see Tables 11 and  12), PACS (employing RLOO) demonstrates highly comparable performance with the Dr.
-GRPO variant.
-This result aligns with expectations, as RLOO and Dr.
-GRPO exhibit only subtle algorithmic differences, both enhancing the foundational GRPO through intra-batch relative comparisons.
-On these moderately challenging tasks, such minor algorithmic distinctions do not translate into substantial performance disparities.
-Specifically, Dr.
-GRPO shows marginal superiority at lower $k$ (kk=1, 2), while PACS exhibits superior performance at higher $k$ values; however, the overall performance of both methods remains remarkably similar.
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+PACS 90.45 96.01 96.96 97.40 97.50 97.50
+   - GRPO 89.55 93.72 95.60 96.79 97.59 98.11
+   - Dr. GRPO 90.04 95.50 96.49 97.15 97.47 97.50
+Table 10: Results of different advantage estimators of Qwen3-4B on AMC23. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
 AIME2024 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-PACS 15.80 19.93 24.39 29.75 36.29 43.90 51.21 56.54 59.78
-    - GRPO 12.75 16.24 19.00 21.59 24.88 28.91 33.22 37.14 39.95
-    - Dr. GRPO 15.59 19.67 23.99 29.04 34.78 40.82 46.76 53.08 59.68
-Table 13: Results of different advantage estimators of Qwen2.5-7B on AIME 2024. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+PACS 55.10 72.91 77.93 80.74 82.25 83.13
+   - GRPO 56.02 67.74 77.01 81.32 83.02 83.33
+   - Dr. GRPO 55.78 71.54 76.28 81.27 82.89 83.32
+Table 11: Results of different advantage estimators of Qwen3-4B on AIME2024. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
 AIME2025 ($\text{pass@}k$)
-Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
-PACS 13.33 18.19 23.73 29.65 36.00 42.88 49.45 54.48 58.22
-    - GRPO 8.39 13.01 17.69 21.78 25.77 29.86 34.34 39.26 43.73
-    - Dr. GRPO 13.93 17.15 20.23 24.05 28.68 34.55 41.07 46.12 49.56
-Table 14: Results of different advantage estimators of Qwen2.5-7B on AIME 2025. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
-The superiority of RLOO is most pronounced on the AIME 24 and AIME 25 (Tables 13 and 14).
-Here, the nuanced algorithmic difference between RLOO and Dr.
-GRPO is amplified into a clear performance divergence, with our PACS algorithm demonstrating unequivocal superiority across all values of $k$.
-We posit that this performance gap is rooted in the distinctive ”leave-one-out” credit assignment of RLOO.
-High-difficulty tasks with sparse rewards and high solution variance demand a robust signal for policy updates.
-RLOO’s mechanism—contrasting each sample against the cohort average—provides exactly this stability.
-The practical consequence is substantial: on the AIME 2025 benchmark, for instance, PACS (58.22%) establishes a lead of almost 9 percentage points over Dr.
-GRPO (49.56%) at pass@256.
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+PACS 45.63 61.00 66.18 70.14 72.91 74.80
+   - GRPO 44.01 59.66 65.92 70.94 75.04 78.96
+   - Dr. GRPO 44.95 60.17 65.90 70.85 75.06 79.03
+Table 12: Results of different advantage estimators of Qwen3-4B on AIME2025. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
+BeyondAIME ($\text{pass@}k$)
+Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
+PACS 27.16 41.66 47.96 53.41 57.91 62.10
+   - GRPO 26.41 39.95 45.73 50.54 54.03 56.30
+   - Dr. GRPO 26.67 40.70 46.87 52.33 56.73 60.36
+Table 13: Results of different advantage estimators of Qwen3-4B on BeyondAIME. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
+Our findings indicate that PACS with RLOO exhibits superior performance, particularly on challenging benchamarks.
+As shown in table 13, on the hardest dataset BeyondAIME, PACS achieves a $\text{pass@}1$ of 27.16% and $\text{pass@}64$ of 62.10%, consistently outperforming Dr.
+GRPO (26.67% / 60.36%) and GRPO (26.41% / 56.30%) across all $k$ values.
+While Dr.
+GRPO demonstrates competitive performance on standard benchmarks like MATH 500 and at higher sample counts on AIME 2025 , PACS maintains a critical edge in the low-sample regime ($\text{pass@}1$) across most tasks (e.g., 90.45% vs.
+90.04% on AMC 23 ).
+We posit that this performance advantage stems from the algorithmic property of RLOO: by contrasting a single sample with the leave-one-out average rather than the group mean including itself, RLOO minimizes the bias in advantage estimation, yielding more precise signals for credit assignment in complex reasoning paths.
+Appendix E Diversity Score Definition
+To rigorously quantify the semantic diversity of the generated reasoning paths, we utilize an embedding-based metric.
+Let $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{N}\}$ denote the evaluation dataset consisting of $N$ queries.
+For each query $q_{i}$, the model generates a set of $K$ candidate outputs, denoted as $\mathcal{O}_{i}=\{o_{i,1},o_{i,2},\dots,o_{i,K}\}$.
+We employ a pre-trained embedding model $E(\cdot)$ (specifically Qwen3-0.6B-embedding (Zhang et al., 2025) in our experiments) to map each textual output $o_{i,j}$ into a $d$-dimensional dense vector representation:
+$\mathbf{v}_{i,j}=E(o_{i,j})\in\mathbb{R}^{d}$ (12)
+For a specific query $q_{i}$, the semantic similarity between two output vectors $\mathbf{v}_{i,m}$ and $\mathbf{v}_{i,n}$ is measured using cosine similarity:
+$\text{sim}(\mathbf{v}_{i,m},\mathbf{v}_{i,n})=\frac{\mathbf{v}_{i,m}\cdot\mathbf{v}_{i,n}}{\|\mathbf{v}_{i,m}\|\|\mathbf{v}_{i,n}\|}$ (13)
+The diversity score for query $q_{i}$, denoted as $Div(q_{i})$, is defined as the average pairwise cosine distance among the generated responses.
+To ensure an unbiased estimate, we compute the average over all unique pairs $(m,n)$ where $1\leq m<n\leq K$.
+Since cosine distance is defined as $1-\text{similarity}$, the formulation is as follows:
+$Div(q_{i})=1-\underbrace{\frac{1}{\binom{K}{2}}\sum_{1\leq m<n\leq K}\text{sim}(\mathbf{v}_{i,m},\mathbf{v}_{i,n})}_{\text{Average Pairwise Similarity}}$ (14)
+Finally, the global diversity metric $S_{div}$ for the model is calculated by aggregating the diversity scores across all queries in the dataset:
+$S_{div}=\frac{1}{N}\sum_{i=1}^{N}Div(q_{i})$ (15)
+A higher $S_{div}$ value indicates greater semantic variation among the sampled responses, reflecting a broader exploration of the solution space.